### v2

# UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models

Abstract
The rapid advancement of diffusion models (DMs) has not only transformed various real-world industries but has also introduced negative societal concerns, including the generation of harmful content, copyright disputes, and the rise of stereotypes and biases. To mitigate these issues, machine unlearning (MU) has emerged as a potential solution, demonstrating its ability to remove undesired generative capabilities of DMs in various applications. However, by examining existing MU evaluation methods, we uncover several key challenges that can result in incomplete, inaccurate, or biased evaluations for MU in DMs. To address them, we enhance the evaluation metrics for MU, including the introduction of an often-overlooked retainability measurement for DMs post-unlearning. Additionally, we introduce UnlearnCanvas, a comprehensive high-resolution stylized image dataset that facilitates us to evaluate the unlearning of artistic painting styles in conjunction with associated image objects. We show that this dataset plays a pivotal role in establishing a standardized and automated evaluation framework for MU techniques on DMs, featuring 7 quantitative metrics to address various aspects of unlearning effectiveness. Through extensive experiments, we benchmark 5 state-of-the-art MU methods, revealing novel insights into their pros and cons, and the underlying unlearning mechanisms. Furthermore, we demonstrate the potential of UnlearnCanvas to benchmark other generative modeling tasks, such as style transfer. The UnlearnCanvas dataset, benchmark, and the codes to reproduce all the results in this work can be found at https://github.com/OPTML-Group/UnlearnCanvas.
Machine Learning
1 Introduction
In recent years, the field of text-to-image technology, epitomized by Diffusion Models (DMs), has seen remarkable progress, demonstrating an unprecedented ability to generate high-resolution, high-quality images across a wide array of subjects (Rombach et al., 2022; Midjourney et al.,, 2023). These models, widely accessible to public, find applications in diverse sectors, including advertising, the creative arts (Fotor et al.,, 2022), and forensic sketching (LabLab AI et al.,, 2023). Central to the training of DMs is their reliance on the data source from the Internet (Schuhmann et al., 2022) to cover a wide range of content. However, this inclusiveness comes with challenges, including potential concerns about harmful generation (Schramowski et al., 2023), copyright issues (Krietzberg, 2023), and biases/stereotypes (Vinker et al., 2023; Turk, 2023).
Figure 1: An illustration of the task MU for DMs. The pretrained model contains the generation capability of different concepts in different domains. MU aims to erase the generation of a certain unlearning target concept without harming the generation of others.
To alleviate the negative social impacts associated with DMs, machine unlearning (MU) techniques are catching increasing attention and have been studied in the field of vision generative models (Schramowski et al., 2023; Gandikota et al., 2023a, b; Kumari et al., 2023; Fan et al., 2023; Heng & Soh, 2023). MU for DMs (see Fig. 1 for an illustration) aims to erase the model’s capacity to generate images related to a specific concept, essentially erasing that concept from the pre-trained model. This concept can be provided in the form of text prompts (i.e., text conditions of DMs), covering a wide range from concrete object entities to abstract notions, such as artistic styles. We refer readers to Sec. 2 for more related work on MU for DMs.
Unlike MU studies on image classification (Izzo et al., 2021; Golatkar et al., 2020; Becker & Liebig, 2022; Thudi et al., 2022; Warnecke et al., 2021; Jia et al., 2023), we lack evaluation benchmarks for MU in DMs. The current primary choice for assessing the performance of unlearned DMs (i.e., DMs post unlearning) is the Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023). This dataset consists of real-world image-to-text prompts designed for erasing concepts like ‘nudity’ and ‘violence’. However, I2P mainly contains isolated unlearning targets and focuses on safety issues related to image generation.
To enhance the assessment of MU in DMs and establish a standardized evaluation framework, we propose to develop a new benchmark dataset, referred to as UnlearnCanvas, which is designed to evaluate the unlearning of artistic painting styles along with associated image objects. Unlike I2P, UnlearnCanvas can be used to assess the removal of copyrighted information, such as artistic style, and it allows for increased diversity by including various object classes and styles. We summarize our contributions below.
$\bullet$ We conduct a systematic review of existing MU methods for DMs and identified three unresolved challenges in their evaluations, such as limited quantitative assessments and a lack of retainability evaluation of unlearned DMs.
$\bullet$ We propose UnlearnCanvas, a large-scale high-resolution stylized image dataset with 24,000 labeled images, 60+ styles, and 20+ object topics. Its style-object dual supervision and high stylistic consistency help us to improve the precision of MU evaluation for DMs.
$\bullet$ We develop a standardized evaluation pipeline for MU in DMs under UnlearnCanvas, which measures unlearning effectiveness, generation retainability, generation quality, and unlearning efficiency.
$\bullet$ We benchmark state-of-the-art MU methods for DMs using UnlearnCanvas, validating the significance of our proposed evaluation metrics and providing insights into the strengths, weaknesses, and unlearning mechanisms.
2 Related Work and Preliminaries
Machine unlearning (MU) for generative models in vision. MU was originally developed to mitigate the (potentially detrimental) influence of specific data points in a pretrained ML model, without necessitating a complete retraining of the model after removing these unlearning data (Cao & Yang, 2015; Bourtoule et al., 2021; Xu et al., 2023). The significance of MU has emerged with the purpose of data privacy protection, in response to regulations such as ‘the right to be forgotten’ (Hoofnagle et al., 2019). However, it has rapidly gained recognition for its role in promoting security, safety, and trustworthiness of ML models. Examples include defense against backdoor attacks (Liu et al., 2022; Jia et al., 2023), fairness enhancement (Chen et al., 2023; Oesterling et al., 2023), and controllable federated learning (Halimi et al., 2022; Jin et al., 2023).
More recently, unlearning has proven crucial in text-to-image generation via DMs to prevent the production of harmful, private, or illegal content (Gandikota et al., 2023a; Zhang et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023; Kumari et al., 2023; Heng & Soh, 2023). For example, DMs have demonstrated a susceptibility to generating NSFW (Not Safe For Work) images when conditioned on inappropriate text prompts (e.g., ‘nudity’ and ‘violence’) (Schramowski et al., 2023; Zhang et al., 2023b; Chin et al., 2023). Similarly, MU has also found applications in preventing the generation of copyrighted information, such as the misuse of artistic painting styles (Gandikota et al., 2023a; Zhang et al., 2023a). In summary, MU can be regarded as a (post-training) model edit operation, aimed at preventing DMs from producing undesired or inappropriate images.
Our setup and objective. Let us denote a pretrained DM by $\bm{\theta}_{\mathrm{o}}$, capable of synthesizing high-quality images conditioned on a provided text prompt $c$ (e.g., ‘A painting of a cat in Van Gogh style’). In this work, we focus on the task of generating images in various painting styles. When there is a request to adapt the model $\bm{\theta}_{\mathrm{o}}$ to prevent the generation of images, e.g., in the ‘Van Gogh’ style, the MU problem arises, with the ‘Van Gogh’ style being the specified unlearning target, also known as the erasing concept (Gandikota et al., 2023a). See Fig. 1 for a visual representation of the concept.
In formulating MU methods, the traditional setup presupposes access to the training dataset $\mathcal{D}$, where $\mathcal{D}_{\mathrm{f}}\subset\mathcal{D}$ represents the forget set that revolves around the unlearning target. A commonly used gold standard of MU is to retrain the model from scratch over the retain set $\mathcal{D}_{\mathrm{r}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$. However, commencing the training process from scratch becomes practically unfeasible because of (1) the substantial computational expense it entails, (2) the difficulty in accurately identifying $\mathcal{D}_{\mathrm{f}}$, and (3) the absence of access to the original training data source.
To overcome the aforementioned difficulties (1)-(3), the current research mainstream (Gandikota et al., 2023a; Kumari et al., 2023; Zhang et al., 2023a; Gandikota et al., 2023b; Heng & Soh, 2023; Fan et al., 2023) has focused on developing approximate yet efficient unlearning methods as practical alternatives to full retraining. The approximation introduced by current unlearning methods raises the question that we aim to address in this work:
How can we accurately assess unlearning performance and establish a benchmark for existing unlearning methods?
3 Unresolved Challenges in MU Evaluation
This work will develop a new benchmark dataset, referred to as UnlearnCanvas (detailed in Sec. 4), focusing on the task of unlearning artistic painting styles in conjunction with associated image objects. With the use of UnlearnCanvas, a text prompt for DMs can be constructed as ‘A painting of a {object} in {style}’. And the unlearning target can be explicitly designated as either ‘style’ or ‘object’. To be specific, UnlearnCanvas is targeted for the open challenges (C1-C3) when building a rigorous quantitative evaluation framework for MU in DMs. These challenges are consolidated through a careful examination of existing studies for object and/or style unlearning in Tab. 1.
Table 1: An overview of the scope and evaluation methods of the existing MU methods for DMs. The symbol denotes neither objects nor styles are covered.
Related work Unlearning targets Quantitative evaluation
Unlearning
effectiveness
Retainability
ESD (Gandikota et al., 2023a)
5 artistic styles
11 objects
objects only objects only
CA (Kumari et al., 2023)
4 artistic styles
4 objects
objects only
FMN (Zhang et al., 2023a)
1 artistic style
7 objects
objects only
UCE (Gandikota et al., 2023b)
7 artistic settings [^1] UCE adopted a multi-concept unlearning setting, where different numbers of styles are unlearned at the same time and total up to 1000 styles are considered.
11 objects
objects only objects only
SA (Heng & Soh, 2023)
1 artistic style
31 objects
objects only
SalUn (Fan et al., 2023)
No artistic styles
11 objects
objects only objects only
(C1) The absence of a diverse yet cohesive unlearning target repository. As shown in Tab. 1, the assessment of MU for DMs, in terms of both effectiveness (concept erasure performance) and retainability (preserved generation quality under non-forgotten concepts), is typically conducted using manually selected targets, often chosen from a limited pool of unlearning targets. A critical aspect missing from MU evaluation is the recognition that unlearning targets related to styles and objects should not be isolated within the dataset used for MU assessment. Instead, they should be considered in a manner that exposes their potential interconnected influences on other targets.
Table 2: Classification accuracy of a style classifier (ViT-L/16) finetuned on WikiArt (Phillips & Mackintosh, 2011). The classifier is tested on the test set and the image set generated by SD v1.5 and v2.0 with the prompt ‘An painting in {artist} style’, where the {artist} is traversed over the WikiArt labels (10 images/prompt).
WikiArt
Train Set
WikiArt
Test Set
Images by
SD v1.5
Images by
SD v2.0
Accuracy 99.8% 85.4% 56.7% 41.2%
(C2) The lack of precision in evaluation. Artistic styles can be challenging to precisely define and distinguish, which makes quantifying their impact on unlearning effectiveness and retainability difficult. To illustrate this challenge, we conducted a preliminary investigation in Tab. 2 using a ViT-based classifier trained on the WikiArt dataset (Phillips & Mackintosh, 2011). This dataset categorizes paintings into 139 distinct styles, each annotated with the creator’s name. Despite training, the ViT-large model achieved only an $84.32\%$ accuracy on the test dataset. This result underscores the inherent complexity of classifying artistic styles, causing inaccuracies in quantitative MU evaluations. The classifier’s suboptimal accuracy on DMs generated images further increases the prediction error (around $50\%$ accuracy). This complicates precise quantitative evaluations.
(C3) The lack of a systematic study on ‘retainability’ of DMs post unlearning. As indicated by the ‘retainability’ column in Tab. 1, assessing the capacity of unlearned DMs to maintain image generation under innocent concepts is notably lacking. However, retainability is a critical indicator of potential ‘side effects’ of an unlearning method. In conjunction with challenge (C1), the inclusion of diverse unlearning targets, encompassing both styles and objects, allows for the evaluation of retainability from various perspectives. For instance, when unlearning an artistic style, it is crucial to consider the retainability of image generation abilities for other styles as well as for objects.
The challenges (C1)-(C3) outlined above have driven us to develop UnlearnCanvas as a solution for the systematic and comprehensive evaluation of MU in DMs. We will show in Sec. 4 that UnlearnCanvas includes a diverse range of artistic styles and objects. Within this dataset, we can explore the interrelationships between paired image styles and image objects to address (C1). Each style category in UnlearnCanvas maintains high intra-style consistency, facilitating accurate style and object classification and enabling post-unlearning quantitative evaluation for (C2). To address (C3), we can evaluate two types of retainability, providing a precise and comprehensive assessment of the potential side effects induced by MU methods.
4 Our Proposal: UnlearnCanvas Dataset
Construction of UnlearnCanvas. As motivated in Sec. 2, UnlearnCanvas is created for ease of MU evaluation in DMs. Its construction involves two main steps: seed image collection and subsequent image stylization; see Fig. 2 for a schematic overview.
Seed image collection. UnlearnCanvas includes images in $60$ unique artistic styles across $20$ distinct objects. The style categories are derived from a set of high-resolution photos, termed seed images. These seed images consist of original real-world photos from Pexels (Pexels et al.,, 2023). In total, there are 400 seed images across 20 object classes.
Image stylization. After collecting the seed images, the next step in dataset development is the stylization process. During this phase, each seed image is transformed into 60 predetermined artistic styles. The goal is to ensure that the resulting images maintain high stylistic consistency while preserving the content details present in the seed images. This stylization process is carried out using the services provided by (Fotor et al.,, 2022). Once all the seed images have been stylized in all 60 styles, the dataset is constructed with super-high-resolution images and organized in a hierarchical structure that balances both style and object categories.
Figure 2: An illustration of the key steps when curating UnlearnCanvas and its key features.
Advantages of UnlearnCanvas. Next, we will highlight the advantages (A1-A3) of UnlearnCanvas to address the MU evaluation challenges (C1-C3) discussed in Sec. 3.
(A1) Style-object dual supervision enables a rich unlearning target bank for comprehensive evaluation. UnlearnCanvas offers a diverse range of unlearning targets by annotating each image with both style and object labels. This is in response to (C1). We have categorized the unlearning targets into two main domains: a style domain comprising $60$ artistic styles, and an object domain including $20$ object classes. This setup allows us to investigate the impact of style unlearning on the fidelity of generated objects in images and vice versa. Besides its role in MU evaluation, UnlearnCanvas could also serve as an annotated testbed for assessing style transfer accuracy and visual in-context learning performance; see details in Appx. C.
(A2) High stylistic consistency ensures precise style definitions and enables accurate quantitative evaluations. In UnlearnCanvas, images labeled with the same style exhibit high intra-style consistency and distinct inter-style differences, as exemplified in Fig. 2. This facilitates the distinction between different styles, effectively addressing challenge (C2). In a similar study to that presented in Tab. 2, we observe that the styles within UnlearnCanvas can be readily classified using ViT-Large (Dosovitskiy et al., 2020), achieving an accuracy of $98\%$ for images generated by SD finetuned on UnlearnCanvas. We refer readers to Tab. A2 for detailed results.
Figure 3: Evaluation metric illustration, with unlearning the ‘Van Gogh style’ as a case study. ✓ and ✗ denote desired and undesired results post unlearning.
(A3) Enabling in-depth retainability analyses for MU evaluation. Using UnlearnCanvas, we can assess DMs’ retainability post-unlearning from two perspectives: in-domain retainability, which evaluates the model’s performance on test-time generations within the same domain as the unlearning target, and cross-domain retainability, which assesses the model’s ability to generate concepts from a different domain. For example, as depicted in Fig. 3, we can define the ‘domain’ as either ‘artistic style’ or ‘object class’. If the unlearning target is the ‘Van Gogh’ style (i.e., the concept to be erased), then generating images in all styles except ‘Van Gogh’ but within the same object class (i.e., ‘Dog’ in Fig. 3) represents the in-domain case. Conversely, out-of-domain refers to generating images with different object classes (e.g., ‘Horses’ in Fig. 3), i.e., objects different from the one associated with the unlearning target.
Figure 4: Image examples with the label “Van Gogh Style” from WikiArt (Phillips & Mackintosh, 2011) and UnlearnCanvas. Images of the same artistic style in UnlearnCanvas exhibit high stylistic consistency compared to WikiArt.
Figure 5: An illustration of the evaluation pipeline proposed in this work using UnlearnCanvas when unlearning a specific target concept ‘Van Gogh Style’. Unlearning performances (including the unlearning effectiveness and retainability) are quantitatively assessed (marked in blue) to accurately reflect the unlearning performance portrait of each method. The unlearning target of the pipeline could traverse all the styles and objects to achieve a complete and comprehensive evaluation.
UnlearnCanvas vs. WikiArt.
To the best of our knowledge, WikiArt (Nichol, 2016) is the most relevant baseline dataset to ours. In Tab. A1, we provide a direct comparison of the key attributes of these two datasets. UnlearnCanvas differs from WikiArt in the following aspects. First, UnlearnCanvas includes a greater number of high-resolution images ($15$M) compared to WikiArt ($2$M), a factor that may enhance the training of state-of-the-art DMs. Second, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness, as illustrated in Fig. 4, where images labeled with ‘Van Gogh Style’ from both datasets are compared. In UnlearnCanvas, the images exhibit high stylistic consistency, while WikiArt lacks the necessary clarity for precise assessment. This will hamper the MU evaluation as discussed in the challenges (C2) and (C3).
5 Evaluation Pipeline
In this section, we introduce the evaluation pipeline and the benchmarked MU methods for UnlearnCanvas. Fig. 5 illustrates the proposed evaluation pipeline for MU in DMs with a designated unlearning target. It comprises four phases (I-IV) to evaluate unlearning effectiveness, retainability, generation quality, and efficiency.
Phase I: Testbed preparation (Fig. 5a). Prior to unlearning, we commence by fine-tuning a specific DM, given by SD (StableDiffusion) v1.5, on UnlearnCanvas for text-to-image generation. We also train a vision transformer (ViT-Large) (Dosovitskiy et al., 2020) for the purposes of style and object classification after unlearning. We ensure that both the SD model and classifiers attain high performance in image generation and classification across all styles and objects. See detailed testbed preparation in Appx. A.
Phase II: Machine unlearning (Fig. 5b). Next, we utilize the selected MU methods for benchmarking to update the DM acquired in Phase I, aiming to eliminate a designated unlearning target (e.g., ‘Van Gogh style’). This results in the unlearned model (denoted as ${\bm{\theta}}_{\mathrm{u}}$), which is the DM updated to avoid generating the unlearning target. This unlearned DM will be used for unlearning performance evaluation.
Phase III: Answer set generation (Fig. 5c). Further, we utilize the unlearned model ${\bm{\theta}}_{\mathrm{u}}$ to generate a set of images conditioned on both the unlearning-related prompts and other innocent prompts. To comprehensively evaluate unlearning performance, three types of answer sets are generated:
Unlearning effectiveness: Images are generated in response to prompts that involve the unlearning target (e.g., any prompt like ‘An image in Van Gogh style’ for the case in Fig. 5). A successfully unlearned model will avoid generating images that match the unlearning target.
In-domain retainability: Images are generated with prompts in the same ‘domain’ as the unlearning target (e.g., style-related prompts for style unlearning), where the prompts can cover all the other styles provided in UnlearnCanvas except the one to be unlearned, e.g., ‘Van Gogh style’ in Fig. 5. This assesses the unlearned model’s ability to retain generation quality within the same domain.
Cross-domain retainability: Images are generated with prompts in other domains. For instance, in the case of style unlearning as shown in Fig. 5, the object domain is considered separate from the style domain. Prompts related to objects are used to evaluate the DM’s cross-domain generation capability after style unlearning.
Table 3: Performance overview of different MU methods evaluated with UnlearnCanvas dataset. The performance assessment includes unlearning accuracy (UA), in-domain retain accuracy (IRA), cross-domain retain accuracy (CRA), and FID. Symbols $\uparrow$ or $\downarrow$ denote whether larger or smaller values represent better performance. Results are averaged over all the style and object unlearning cases. The best performance regarding each metric is highlighted in bold.
Method Effectiveness Efficiency
Style Unlearning Object Unlearning FID ($\downarrow$) Time Memory Storage
UA ($\uparrow$) IRA ($\uparrow$) CRA ($\uparrow$) UA ($\uparrow$) IRA ($\uparrow$) CRA ($\uparrow$) (s) ($\downarrow$) (GB) ($\downarrow$) (GB) ($\downarrow$)
ESD (Gandikota et al., 2023a) $\mathbf{98.58}\%$ $80.97\%$ $93.96\%$ $92.15\%$ $55.78\%$ $44.23\%$ $65.55$ 6163 17.8 4.3
FMN (Zhang et al., 2023a) $88.48\%$ $56.77\%$ $46.60\%$ $45.64\%$ $90.63\%$ $73.46\%$ $131.37$ $\mathbf{350}$ 17.9 4.2
UCE (Gandikota et al., 2023b) $98.40\%$ $60.22\%$ $47.71\%$ $\mathbf{94.31}\%$ $39.35\%$ $34.67\%$ $182.01$ 434 $\mathbf{5.1}$ $\mathbf{1.7}$
CA (Kumari et al., 2023) $60.82\%$ $\mathbf{96.01}\%$ $92.70\%$ $46.67\%$ $90.11\%$ $81.97\%$ $\mathbf{54.21}$ 734 10.1 4.2
SalUn (Fan et al., 2023) $86.26\%$ $90.39\%$ $\mathbf{95.08}\%$ $86.91\%$ $\mathbf{96.35}\%$ $\mathbf{99.59}\%$ $61.05$ 667 30.8 4.0
Phase IV: MU performance assessment (Fig. 5d). After Phase III, the answer set undergoes style/object classification for unlearning performance assessment. This classification results in three quantitative metrics ①-③.
Unlearning accuracy (UA): It represents the proportion of images generated by the unlearned DM using the unlearning target-related prompt (e.g., the Van Gogh style prompt in Fig. 5), which are not correctly classified into the corresponding class for the case. Therefore, a higher UA indicates better unlearning performance in preventing image generation from the unlearning target-related prompts.
In-domain retain accuracy (IRA): It is given by the classification accuracy of images generated by the unlearned DM using innocent prompts (not relevant to the unlearning target) in the same domain, e.g., all styles other than the Van Gogh style in Fig. 5.
Cross-domain retain accuracy (CRA): Similar to IRA, it is given by the classification accuracy of images generated by the unlearned DM using innocent prompts but in different domains, e.g., all object categories in Fig. 5.
In addition to the above accuracy metrics, we also utilize the FID score (④) to evaluate the distribution-wise generation quality of the unlearned DM. Furthermore, we will also monitor the efficiency of the unlearning process, taking into account factors such as run-time (⑤), storage space requirements (⑥), and memory costs (⑦). In total, the performance evaluation involves 7 metrics (①-⑦).
MU methods for DMs to be benchmarked. In this work, we assess the 5 most recently proposed MU methods for DMs, including ESD (Gandikota et al., 2023a), FMN (Zhang et al., 2023a), UCE (Gandikota et al., 2023b), CA (Kumari et al., 2023), and Salun (Fan et al., 2023). We remark that SA (Heng & Soh, 2023), a recently proposed method, is excluded from our evaluation due to its excessive computational resource requirements and time consumption. Detailed training settings for each unlearning method can be found in Appx. A.
Figure 6: Performance visualization of Table 3. For UA, IRA, and CRA, the results are averaged over all the style and object unlearning scenarios. Other metrics undertake the inverse operation as a smaller values represent better performance. Results are normalized to $0\%\sim 100\%$ per metric.
Figure 7: Left: Heatmap visualization of the unlearning and retain accuracy of ESD on UnlearnCanvas. The $x$-axis shows the tested concepts for image generation using the unlearned model, while the $y$-axis indicates the unlearning target. Concept types are distinguished by color: styles in blue and objects in orange. The figure is separated into different regions to represent corresponding evaluation metrics and the unlearning scopes ($A$ for UA, $B$ for IRA, $C$ for CRA; ‘1’ for style unlearning, ‘2’ for object unlearning). Diagonal regions ($A1$ and $A2$) indicate unlearning accuracy, and off-diagonal values ($B1$, $B2$, $C1$, and $C2$) represent retain accuracy. Higher values in lighter color denote better performance. The first row serves as a reference for comparison before unlearning. Zooming into the figure is recommended for detailed observation. Right: Representative cases illustrating each region with images generated before and after unlearning a specific concept.
6 Experiment Results
In this section, extensive experiments are conducted on UnlearnCanvas following the evaluation pipeline in Sec. 5. We show that previous MU evaluation, primarily focused on unlearning effectiveness, may present a biased view without considering retainability assessment (Tab. 3). We also find that preserving cross-domain retainability, as quantified by CRA, is more challenging than in-domain retainability (measured by IRA), a previously overlooked aspect in the literature. Furthermore, we show that no single MU method consistently excels across all metrics, emphasizing the potential for advancements in existing methods (Fig. 6). In addition, we present a detailed case study on ESD, showcasing how our benchmark offers a comprehensive analysis of an unlearning method’s mechanism on forgetting and retaining information (Fig. 7). Last but not least, we show that different MU methods display distinct unlearning characteristics, which are inherently linked to their underlying unlearning mechanisms (Fig. 8).
Overall performance assessment of MU for DMs. In Tab. 3, we provide an overview of the performance of existing unlearning methods, using the metrics discussed in Sec. 5. The key results are summarized below.
(1) Retainability is essential for comprehensive assessment of MU in DMs. Previous evaluations heavily relied on UA (unlearning accuracy), but our assessment reveals a crucial insight: relying solely on UA can provide a skewed view of MU method performance. For example, when unlearning artistic painting styles (that we call style unlearning), ESD and UCE achieve similar UA (both around $98\%$), but their retainability (IRA and CRA) differs significantly, with one over $80\%$ and the other around $60\%$. Similar conclusions apply to unlearning an image object (that we call object unlearning). These findings emphasize the importance of including retainability metrics in the evaluation framework, as demonstrated by UnlearnCanvas.
(2) CRA is harder to retain than IRA. In both style and object unlearning, CRA presents a greater challenge for MU methods compared to IRA. This trend is consistent across various methods. For example, the UCE method (Gandikota et al., 2023b) exhibits a significant gap between its IRA and CRA, both in style ($60.22\%$ vs. $47.71\%$) and object ($39.35\%$ vs. $34.67\%$) unlearning scenarios. Similar patterns are observed with other methods like FMN and CA. This pattern suggests that while MU methods are somewhat effective at preserving concepts within the same domain (in terms of IRA), they face greater challenges in maintaining performance on unlearning target-unrelated concepts across domains (CRA), which has been overlooked in previous studies due to the absence of a multi-label dataset like UnlearnCanvas for systematic unlearning analysis.
(3) One MU method can perform differently in different domains. Take the FMN method (Gandikota et al., 2023b) as an example. It exhibits an evident performance difference between the two domains, achieving a UA of $88.48\%$ in style unlearning but a much lower UA of $45.64\%$ in object unlearning. In addition, its IRA and CRA metrics show a notable gap, with an IRA of $90.63\%$ for object unlearning compared to $56.77\%$ for style unlearning. These examples highlight that a method’s effectiveness in unlearning is not consistent across different scopes.
(4) No single method can excel in all aspects. Each method exhibits unique strengths and weaknesses. For example, while ESD (Gandikota et al., 2023a) typically leads an improvement in UA, it may lag behind in terms of IRA and CRA, e.g., in the case of object unlearning. Also, the SalUn method (Fan et al., 2023) shows exceptional retainability, with a CRA of $99.59\%$ in object unlearning, but its UA in the same category is lower at $86.91\%$. Furthermore, the UCE method (Kumari et al., 2023) enjoys remarkably high UA and efficiency but suffers from degradation in FID and retainability. For better visualization, in Fig. 6, we show this pattern by plotting the performance of different metrics reported in Tab. 1, illustrating that no method dominates across all radar chart dimensions. This underscores the complexity of unlearning in generative models and the need for further advancement.
Refer to caption Refer to caption Refer to caption
(a) ESD (b) SalUn (c) Illustration
Figure 8: Visualization of the unlearning directions of (a) ESD and (b) SalUn. This figure illustrates the conceptual shift of the generated images of an unlearned model conditioned on the unlearning target. Images generated by the post-unlearning models are classified and used to understand this shift. Edges leading from the object in the left column to the right signify that images generated conditioned on unlearning targets are instead classified as the shifted concepts after unlearning. This reveals the primary unlearning direction for each unlearning method. The most dominant unlearning direction for an object is visualized. Figure (c) provides visualizations of generated images using the prompt template ‘A painting of {object} in Sketch style.’ with object being each unlearning target.
A closer look into retainability: A case study on ESD. We next perform a more in-depth analysis of DMs’ retainability after unlearning, with a focus on the ESD method (Gandikota et al., 2023a), which has shown superiority in Tab. 3 for style unlearning. In Fig. 7 (left), we present a heatmap that illustrates the per-style/object unlearning accuracy (diagonal) and retain accuracy (off-diagonal) for ESD. The heatmap segments are labeled to indicate different regions of interest (ROIs) related to our evaluation metrics: A1 and A2 for Unlearning Accuracy (UA) in style and object unlearning, B1 and B2 for In-Domain Retain Accuracy (IRA) in style and object unlearning, and C1 and C2 for Cross-Domain Retain Accuracy (CRA) in style and object unlearning. We highlight the following observations.
First, we observe distinct behaviors of ESD in unlearning styles compared to objects. The ROI C2 appears darker than C1, suggesting a reduced ability to retain styles during the unlearning of objects. Similar patterns are observable between B2 and B1. These findings indicate that an MU method might demonstrate a preference within a specific domain but face challenges in others. For a clearer explanation, Fig. 7 (right) provides image generation examples of the DM before and after unlearning within different ROIs.
Second, we observe that style/object unlearning is relatively easier compared to retaining the generation performance of unlearned DMs conditioned on unlearning-unrelated prompts. This is evident in the contrast between ROIs A1/A2 and ROIs B1/B2 or C1/C2. One exception is the case of unlearning the object ‘Sea’, which exhibits relatively low UA in A2. This may be attributed to the image generation post-unlearning still resembling sea waves, as shown by the image examples in Fig. 7 (right) for A2.
Third, we present a similar accuracy heatmap for SalUn in Fig. A1. In comparison to ESD, SalUn demonstrates more consistent performance across different unlearning scenarios. However, SalUn does not achieve the same level of UA as ESD, as evidenced by the lighter diagonal values in Fig. A1. This observation underscores the trade-off between unlearning effectiveness and retainability of MU in DMs.
Understanding unlearning method’s behavior via unlearning directions. As noted earlier, different unlearning methods display distinct unlearning behaviors. To gain insights into the underlying reasons for these differences, Fig. 8 (a) and (b) visualize the ‘unlearning directions’ for ESD and SalUn, respectively. These unlearning directions are determined by connecting the unlearning target with the predicted label of the generated image from the unlearned DM conditioned on the unlearning target. As shown in Fig. 8 (a), ESD demonstrates a focused shift in image generation after object unlearning, with a predominant transition towards generating images labeled by ‘Sea’ and ‘Trees’. This behavior arises from ESD’s optimization process, designed to steer the generation of the DM away from a predefined concept. Consequently, images generated by the ESD-induced unlearned model consistently lack clearly identifiable objects, resembling waves and trees, which leads to their classification into the ‘Sea’ and ‘Trees’ classes; see Fig. 8 (c) for examples of generated images. In contrast, SalUn exhibits a more diverse range of unlearning directions, shifting images to 11 different objects. This diversity results from SalUn’s requirement to replace the unlearning target with a random concept. As shown in Fig. 8 (c), images generated by SalUn post-object unlearning still maintain some object contours (different from the original unlearning target) and better retain style information compared to ESD.
Expanding applications to style transfer. In addition to evaluating DMs’ unlearning performance, it’s worth noting that UnlearnCanvas has the potential to extend its utility to another generative modeling task: style transfer (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021). One significant challenge in evaluating style transfer methods is the absence of ground truth images for the given reference style image and content image (Cai et al., 2023). Thus, quantitative evaluations are also very limited in the literature. In Appx. C, we utilize UnlearnCanvas to assess 9 latest style transfer methods, providing enriched and precise evaluations. While not the primary focus of this work, this illustrates the great potential of UnlearnCanvas to contribute broader visual generative modeling tasks.
7 Conclusion
In this paper, we systematically reviewed the existing the stateful MU (machine unlearning) methods on DMs (diffusion models), and pinpointed the key challenges rooted in their evaluation systems, which could lead to incomplete, inaccurate, and biased assessments. In response, we propose UnlearnCanvas, a high-resolution stylized image dataset facilitating to build a comprehensive evaluation system for MU methods, together with a novel systematic retainability measurements, that have been long overlooked but are essential in a all-round evaluation. By benchmarking 5 stateful MU methods, we reveal novel insights into the their strengths and weaknesses, deepen the understanding on the underlying mechanism, and further illuminate several promising research directions. In the end, with a case study on style transfer, we show UnlearnCanvas enjoys high versatility and great potential to benchmark other generative modeling tasks beyond MU.
Impact Statement
This work helps improve the assessment and further promotes the advancement of MU (machine unlearning) methods for DMs (diffusion models), which are known to be effective in relieving or mitigating the various negative societal influences brought by the prevalent usage of DMs, which include but are not limited to the following aspects.
$\bullet$ Avoiding Copyright Issues. There is an urgent need for the generative model providers to scrub the influence of certain data on an already-trained model. In January 2023, a notable lawsuit targeted two leading AI art generators, Stable Diffusion (Rombach et al., 2022) and Midjourney (Midjourney et al.,, 2023), for alleged copyright infringement. Concurrently, incidents with the recently released Midjourney V6 (Midjourney et al.,, 2023) also highlighted a visual plagiarism issue on famous film scenes. These instances illuminate the broad copyright challenges inherent in the way of training data collection method of those foundation generative models’ training datasets. MU methods can be used as an effective method to remove the influence of the private data and avoid unnecessary retraining.
$\bullet$ Mitigating biases and stereotypes. Generative AI systems are known to have tendencies towards bias, stereotypes, and reductionism, when it comes to gender, race and national identities (Turk, 2023). For example, a recent study on the images generated with Midjourney revealed, that images associated with higher-paying job titles featured people with lighter skin tones, and that results for most professional roles were male-dominated (Vinker et al., 2023). MU is known to be effective in eliminating biases rooted in the training data. Moreover, UnlearnCanvas offers a flexible framework to benchmark MU techniques against bias removal, allowing for the creation and quantitative control of biases across different object classes for comprehensive bias removal studies.
Appendix
A Detailed Experiment Settings
A.1 Fine-tuning Style and Object Classifier with UnlearnCanvas.
Style and object classifiers need to be trained as part of the test bed proposed in our evaluation pipeline (Fig. 5). Here, we adopted a pretrained ViT-L/16 model (Dosovitskiy et al., 2020) pretrained on ImageNet and fine-tune it on UnlearnCanvas. UnlearnCanvas are split into the train set and test set with a ratio of $9:1$. After hyper-parameter tuning, the classifiers are trained with Adam optimizer with a learning rate of $0.01$ for 10 epochs. We noticed the model training accuracy can readily converge to $100\%$ even with 10 epochs of training.
A.2 Fine-tuning Stable Diffusion Model with UnlearnCanvas.
The other part of the testbed is a diffusion model capable of generating high quality images in all the styles associated with all the objects encompassed in UnlearnCanvas in order to guarantee a trustworthy and unbiased evaluation.
Training settings.
Practically, we fine-tune the pretrained StableDiffusion v1.5 on UnlearnCanvas for 20k steps with a learning rate of $61e-$ 6. Unless otherwise stated, we strictly follow the training configurations used in Stable Diffusion (Rombach et al., 2022). For each image in UnlearnCanvas, we annotate the data with text prompt An image of {object} in {style} style, where the object and style are the corresponding object and style label. For seed images, the style label we use is ‘photo’ style. We use the training scripts provided by Diffuser official tutorial (StableDiffusion) and the pretraining model card is runwayml/stable-diffusion-v1-5. During training, the checkpoints will be saved every 1000 steps.
Evaluation.
To evaluate the quality of the saved checkpoints and selcect the best one for unlearning study, the checkpoints are first used to generate an image set with the same prompt as training (An image of {object} in {style} style) by traversing all the possible style and object labels. Each prompt are used to generate $5$ images with different random seed. The image set for each checkpoint are fed into the style and object classifier trained in Appx. A.1. The model with the highest average performance on all the styles and objects are selected as the testbed for MU study. The classification performance are also used as a reference for later IRA/CRA comparison, which are disclosed in the first row of Fig. 7 (left). In Sec. 6, the experiments were built on the $50$ out of the $60$ styles, which served as the first edition of the dataset, and these styles are listed in Fig. 7. We remark that similar experiments can be done on all the 60 styles.
A.3 Implementation of the Tested Machine Unlearning Methods
In this work, we inspected 5 stateful MU methods for DMs. For each method, we use their publicly released source codes as code bases, which are listed below:
In particular, we adopt the following training settings to adapt the methods to our dataset:
  • ESD: based on the suggestions from the authors, ESD is used to only fine-tune the cross attention-related model weights (ESD-x). Other settings strict follow the ones used in the paper.
  • CA: in order to ablate concepts using CA, we first use ChatGPT to generate a list of simple prompts for each concept (including the styles and the objects), namely anchor prompts. Each anchor prompt is a simple one-sentence description of the unlearning target.
  • UCE: This method requires a guided concept (prompt) for each unlearning concept. For style unlearning, we use the prompt An image in {style*} style as the guided concept, where style* represents the next style in UnlearnCanvas in alphabetical order. Similarly, An image of object* is used for object unlearning, where object* is the next object in UnlearnCanvas in alphabetical order.
  • FMN: This method requires the images associated with the unlearning target. For simplicity and best performance, we randomly select 20 images associated with the unlearning concept. For the first stage of FMN, we run text inversion for 500 steps with a learning rate of $41e-$ 4, and for the second step, we used the inversed text to unlearn the cross attention layers of the model for $100$ steps. The hyper-parameters of learning rate, maximum steps, and tunable parameters (cross-attention or non-cross-attention) are carefully tuned with grid search.
  • SalUn: This method involves two steps, the mask finding (weight saliancy analysis) and the model unlearning. For mask finding, we tuned the mask ratio, while for unlearning, we tune the hyper-parameter learning rate and unlearning intensity. All the parameters are tuned with grid search. For both steps, the mask or model is trained with $10$ epochs.
B Additional Machine Unlearning Experiment Results
Table A1: Comparison with WikiArt, the most relevant dataset containing stylized images to ours. UnlearnCanvas stands out notably from WikiArt due to its characteristics of being supervised, balanced, and maintaining high stylistic cohesiveness.
Dataset
Resolution
(Pixels/Image)
Style-wise
Supervised
High Stylistic
Consistency
Class-wise
Balanced
WikiArt (Nichol, 2016) $\sim 2M$
Style-wise
Object-wise
UnlearnCanvas $\sim 15M$
Style-wise
Object-wise
Comparison to the most relevant dataset WikiArt.
We first make a comprehensive comparison between UnlearnCanvas and WikiArt on their key properties in Tab. A1. As we can see, UnlearnCanvas out performs WikiArt from the several perspectives. First, UnlearnCanvas contain images of much higher resolution (15M vs. 2M). This in particular benefits the training associated with DMs, where larger images in general help the model captures more details, such as the stylistic strokes. Second, the images in UnlearnCanvas are style-wise supervised. This means, for each seed image, a stylized counterpart can be find in each style class. This is beneficial for other tasks than MU, such as style transfer, which can provide a ground truth image for precise and robust evaluation. This will be detailed in Appx. C. Third, UnlearnCanvas enjoys high intra-style stylistic consistency and inter-style stylistic distinctiveness. This means the images of the same label share much more stylistic similarity, which is easier for a deep learning model to learn and to classify. This benefits can also be reflected by the training performance Appx. A.1 using UnlearnCanvas and WikiArt. The results are reported in Tab. 2 and Tab. A2, respectively. As we can see, the classifier is much more easily trained on UnlearnCanvas, justifying the higher discernible features within each style in UnlearnCanvas.
Table A2: Style classification results of a ViT-Large (Dosovitskiy et al., 2020) as a style classifier trained on UnlearnCanvas. After convergence, the classifier is tested on the test set and the image set generated by SD v1.5 finetuned on UnlearnCanvas.
UnlearnCanvas
Train Set
UnlearnCanvas
Test Set
Images by SD v1.5
tuned on UnlearnCanvas
Accuracy $100.0\%$ $99.9\%$ $98.8\%$
A fine-grained comparison per unlearning target: ESD vs. Salun.
Following the analysis of ESD, we next turn our focus to a comparative analysis with Salun, a method that demonstrated a better balance between unlearning and retaining according to Table 3. A similar accuracy heatmap for Salun is presented in Figure A1. Compared to ESD, Salun exhibits more consistent performance across various unlearning scenarios, as indicated by the more uniform color distribution in the heatmap. This also suggests enhanced retainability. However, it is noticeable that Salun does not reach the same level of UA (Unlearning Accuracy) as ESD, as evidenced by the darker diagonal values in Figure A1. This observation reinforces the existence of a trade-off between unlearning effectiveness and retainability in the visual generative MU task, a phenomenon paralleled in other tasks such as classification.
Figure A1: Heatmap visualization of SalUn. The plot setting is identical to Figure 7.
C Broader Use Cases of UnlearnCanvas
Although UnlearnCanvas is originally designed for benchmarking MU methods, we would like to demonstrate its broader use cases of benchmarking more generative modeling tasks, thanks to its good properties discussed in Sec. 4. In this section, we start with a case study on the task of style transfer, which is a much more well-studied topic than MU, but surprisingly also faces great challenges in building up a comprehensive and precise evaluation framework. In the next, we will first dissect the key challenges of the current style transfer evaluation framework, and then demonstrate how UnlearnCanvas efficiently resolves these challenges and further proposes a comprehensive and automated benchmark. Through extensive experiments, we draw demonstrate new insights from these results and illuminate the challenges of the future research directions. In the end, we will discuss the possibility of using UnlearnCanvas to benchmark more generative modeling tasks.
C.1 Benchmarking Style Transfer using UnlearnCanvas
Figure A2: An illustration of the task of style transfer.
Style transfer.
Style transfer is a long-standing topic and focuses on transferring the artistic style from one style image (also known as the reference image) $\mathbf{x}_{s}$ to a target content image $\mathbf{x}_{c}$. The most recent methods typically employ a neural network, denoted by ${\bm{\theta}}_{s}$, to extract style features and perform stylization in a single inference step, expressed as $\hat{\mathbf{x}}_{\mathrm{o}}=f_{{\bm{\theta}}_{s}}(\mathbf{x}_{s},\mathbf{x}_{c})$. Current state-of-the-art (SOTA) style transfer techniques exhibit remarkable generalization capabilities, successfully transferring styles not encountered during training and not requiring any further back-propagations. Figure 1 illustrates the pipeline of this task. Most existing literature (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021) utilize WikiArt (Phillips & Mackintosh, 2011) to provide different styles for training the stylization network ${\bm{\theta}}_{s}$. During the evaluation, the validation set of WikiArt will serve as the style (reference) images, together with the content images from the COCO dataset (Lin et al., 2014), to form a test bed for style transfer and style learning methods. However, such a evaluation scheme has some inherent limitations, which will be detailed below.
Issues and challenges in the evaluating methods for style transfer
Although the task of style transfer has been widely studied, its evaluations are still based on very limited quantitative or even sorely qualitative assessments, potentially leading to incomplete and inaccurate assessments (Cai et al., 2023). Upon examining the evaluation pipelines of over $10$ state-of-the-art (SOTA) style transfer methods, three significant challenges are identified within the current widely accepted evaluation frameworks. $\bullet$ Challenge I (C1): The lack of the ground truth images for style similarity evaluation. Unlike other vision tasks like classification (Singh & Singh, 2020), detection (Zou et al., 2023), and segmentation (Minaee et al., 2021), one of the key shortcomings of the current style transfer evaluation lies in the lack of the ground truth images $\mathbf{x}_{g}$ for the given reference style image $\mathbf{x}_{s}$ and the content image $\mathbf{x}_{c}$. Consequently, existing evaluation metrics, such as the style loss $\ell_{\mathrm{style}}$ (Gatys et al., 2015a, b), has to be calculated with the reference style image $\mathbf{x}_{s}$ as the ground truth $\mathbf{x}_{g}$, namely $\ell_{\mathrm{style}}(\mathbf{x}_{s},\hat{\mathbf{x}}_{o})$, rather than directly using the ground truth $\ell_{\mathrm{style}}(\mathbf{x}_{o},{\mathbf{x}}_{g})$. Obviously, such a indirect evaluation may lead to inaccurate results due to the different contents held in $\mathbf{x}_{s}$ and $\mathbf{x}_{o}$. Existing work has demonstrated that such evaluation metrics can lead to very different conclusion from that of the user study (An et al., 2021). Therefore, the creation of a supervised dataset with ground truth stylized images for every content image under each style is a timely remedy. $\bullet$ Challenge II (C2): The lack of algorithm stability evaluation against varied reference images $\mathbf{x}_{s}$. The evaluation of algorithm stability in style transfer has often been neglected. This assessment requires consistent performance of a method on a content image $\mathbf{x}_{c}$ across different style reference images ${\mathbf{x}_{s}}$ representing the same target artistic style. Ideally, an algorithm should maintain uniform quality across various references within a style, avoiding significant performance variations. Current challenges in such evaluations stem from the lack of reference sets that exhibit high stylistic consistency within each style. Figure 4 showcases examples from the widely-used WikiArt dataset. Despite sharing the same artistic label, images within a row show considerable divergence in visual appearance and style. Consequently, using these images as style references can lead to stable algorithms producing stylistically varied outputs, leading to misleading assessment results. Therefore, creating a dataset with high stylistic uniformity within each style category and clear differentiation between styles is crucial for accurate measurement of algorithm stability.
Refer to caption Refer to caption Refer to caption (a) Style Authenticity (b) Content Integrity (c) Algorithm Stability
Figure A3: An illustration of comprehensive performance metrics for style transfer tasks. (a) Style authenticity measures the style similarity between the generated image and the style image. (b) Content integrity measures the content preservation between the generated image and the content image. (c) Algorithm stability reflects the sensitivity of the algorithm to different style images. For all the metrics in the illustration, “Method II” is always better than “Method I”.
Building up a comprehensive evaluation pipeline with UnlearnCanvas for style transfer.
To revolutionize the evaluation framework with UnlearnCanvas, it is crucial to first understand the components that form a comprehensive evaluation pipeline, ensuring a thorough and impartial assessment of performance. In the realm of style transfer, three pivotal metrics stand paramount, as depicted in Figure,A3:
❶ Style Authenticity: Assesses the extent to which the style of the generated image aligns with that of the provided reference style image(s). ❷ Content Integrity: Measures how well the content features are preserved post style transfer. ❸ Algorithm Stability: Evaluates the algorithm’s robustness against variations in content subjects, target styles, or style reference image selections, while keeping other parameters constant.
In addressing the challenges (C1-C2), UnlearnCanvas proves to be inherently advantageous. First, the style-specific supervision embedded in UnlearnCanvas enables the provision of ground truth for quantitative evaluations of stylized images. Second, the extensive image collection within the same style category in UnlearnCanvas allows for the assessment of algorithm stability by applying style transfer methods to varied reference images. To encompass the aforementioned aspects of style transfer performance, we propose the following quantitative evaluation metrics:
① Style Loss: Utilizes a feature map-based style loss (Deng et al., 2022) to quantify the stylistic dissimilarity between image pairs, effectively representing the inverse of style authenticity. ② Content Loss: Employs a VGG-based, feature-map content loss (Deng et al., 2022; Zhang et al., 2022) to measure the visual dissimilarity between the reference and generated images, essentially mirroring content integrity. ③ Averaged Standard Deviation (STD): Computes the average STD of style and content loss w.r.t. the same reference image, reflecting algorithm stability.
Furthermore, similar to the MU task, we also consider efficiency metrics for each method, including:
④ Average Time Consumption: Measures the time required for performing style transfer. ⑤ Peak GPU Memory Consumption: Records the maximum GPU memory usage during the style transfer process. ⑥ Model Storage Memory Consumption: Assesses the memory requirement for storing the style transfer model.
With the introduction of evaluation metrics (①-⑥), we establish a comprehensive evaluation pipeline for style transfer. The process is as follows:
For each style transfer method, style transfer is executed within each object class. Specifically, for every style, images indexed from $1$ to $18$ serve as style reference images, while seed images corresponding to indices $19$ and $20$ are used as content images. Style and content loss are computed for each pair of style reference and content images, with the stylized images derived from the used seed content images serving as the ground truth for style loss. This results in a total of $60\times 20\times 18\times 2$ experimental trials for each method. For a specific style and content image pair, $18$ experiments are performed to calculate the Standard Deviation (STD) values for both style and content loss. These STD values are then averaged over all content images (amounting to $60\times 20\times 2$ cases). The findings from these comprehensive evaluations are presented in Table A3.
Following this evaluation pipeline, we scrutinized $9$ prominent style transfer methods, including SaNet (Park & Lee, 2019), Mcc (Deng et al., 2021), Mast (Huo et al., 2021), ArtFlow with its two variants (AF-AdaIn and AF-WCT) (An et al., 2021), IE-Contrast (Chen et al., 2021), Cast (Zhang et al., 2022), StyTr2 (Deng et al., 2022), and BLIP (Li et al., 2023).
Table A3: Performance overview of different style transfer methods evaluated with UnlearnCanvas dataset. The performance are assessed from the perspectives of stylistic authenticity (style loss), content integrity (content loss), algorithm stability (standard deviations from different dimensions), and efficiency. For all the metrics, smaller values are always preferred for better performance. The best performance per each metric is highlighted in bold. The standard deviations are first calculated with respect to different styles, object classes, or tested content images and then averaged in order to depict the algorithm stability from different perspectives.
Method Style Loss Content Loss Efficiency
Mean STD (Averaged over) Mean STD (Averaged over) Time (s/image) Memory (GB) Storage (GB)
Style Object Content Image Style Object Content Image
SaNet 23.48 2.73 2.87 1.87 0.85 0.12 0.17 0.09 0.29 2.3 0.11
Mcc 17.92 4.59 4.82 2.14 0.96 0.14 0.21 0.07 0.38 5.4 0.10
Mast 24.10 2.87 3.16 1.74 1.42 0.33 0.34 0.18 2.86 4.8 0.16
AF-AdaIn 20.78 2.96 3.13 1.65 1.09 0.11 0.19 0.05 0.53 6.3 0.08
AF-WCT 20.22 2.94 3.19 1.75 1.02 0.12 0.18 0.05 0.53 6.3 0.08
IE-Contrast 21.27 3.01 3.32 2.05 1.08 0.29 0.31 0.15 0.05 3.8 0.11
Cast 24.01 2.78 2.90 1.35 1.38 0.32 0.40 0.16 0.32 6.7 0.19
StyTr2 19.75 3.04 3.30 1.91 0.62 0.10 0.12 0.04 0.58 3.9 0.21
BLIP 25.43 2.90 3.06 2.03 1.61 0.30 0.34 0.16 8.87 7.2 7.23
Experiment results analysis.
Tab. A3 provides a systematic evaluation of the performance of various methods tested. From the analysis, we can derive several crucial insights:
First, it is evident that no single method excels across all evaluation metrics. Notably, Mcc demonstrates superior performance in maintaining stylistic authenticity, as indicated by a low style loss. Conversely, StyTr2 stands out in preserving content integrity, reflected by its minimal content loss.
Second, the assessment of standard deviation is indispensable for a comprehensive evaluation. The method with the optimal performance does not necessarily exhibit the greatest stability. This is particularly apparent in the style loss evaluation, where Mcc, despite achieving the best result in terms of style loss, exhibits the least stability, denoted by the highest standard deviation.
C.2 Broader Use Cases of UnlearnCanvas
In the preceding section, we demonstrated the application of UnlearnCanvas in refining evaluation metrics and frameworks for style transfer. Beyond this, we recognize the potential of UnlearnCanvas in diverse domains. Here, we delve into two illustrative examples:
Bias mitigation. Bias mitigation in DMs, which are now gaining popularity, can also benefit from UnlearnCanvas. Its hierarchical and balanced architecture enables the deliberate introduction of artificial biases by selectively omitting data from specific groups. For instance, by predominantly excluding images from styles other than the ‘Van Gogh Style’ within the ‘Dogs’ class, DMs fine-tuned on this dataset will inherently exhibit a tendency to generate images of dogs in the Van Gogh style, particularly when the style is not explicitly specified in the prompt. This approach not only allows for the manipulation and quantification of biases but also paves the way for UnlearnCanvas to become a standardized benchmark for bias mitigation, similar to the role of MU for DMs.
Vision in-context learning (V-ICL). V-ICL (Wang et al., 2023b, a; Bai et al., 2023; Wang et al., 2024) is another domain where UnlearnCanvas can be effectively applied. The field of V-ICL is in urgent need of robust, comprehensive methods for the fair assessment of existing models. In this context, the image pairs from UnlearnCanvas are ideally suited for evaluating various tasks such as style transfer, image inpainting, and image segmentation, offering a rich resource for nuanced and quantitative analyses.

### v4

@@
-UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models
+UnlearnCanvas: A Stylized Image Dataset for Enhanced Machine Unlearning Evaluation in Diffusion Models
 Abstract
-The rapid advancement of diffusion models (DMs) has not only transformed various real-world industries but has also introduced negative societal concerns, including the generation of harmful content, copyright disputes, and the rise of stereotypes and biases.
-To mitigate these issues, machine unlearning (MU) has emerged as a potential solution, demonstrating its ability to remove undesired generative capabilities of DMs in various applications.
-However, by examining existing MU evaluation methods, we uncover several key challenges that can result in incomplete, inaccurate, or biased evaluations for MU in DMs.
-To address them, we enhance the evaluation metrics for MU, including the introduction of an often-overlooked retainability measurement for DMs post-unlearning.
-Additionally, we introduce UnlearnCanvas, a comprehensive high-resolution stylized image dataset that facilitates us to evaluate the unlearning of artistic painting styles in conjunction with associated image objects.
-We show that this dataset plays a pivotal role in establishing a standardized and automated evaluation framework for MU techniques on DMs, featuring 7 quantitative metrics to address various aspects of unlearning effectiveness.
-Through extensive experiments, we benchmark 5 state-of-the-art MU methods, revealing novel insights into their pros and cons, and the underlying unlearning mechanisms.
-Furthermore, we demonstrate the potential of UnlearnCanvas to benchmark other generative modeling tasks, such as style transfer.
-The UnlearnCanvas dataset, benchmark, and the codes to reproduce all the results in this work can be found at https://github.com/OPTML-Group/UnlearnCanvas.
-Machine Learning
+The technological advancements in diffusion models (DMs) have demonstrated unprecedented capabilities in text-to-image generation and are widely used in diverse applications.
+However, they have also raised significant societal concerns, such as the generation of harmful content and copyright disputes.
+Machine unlearning (MU) has emerged as a promising solution, capable of removing undesired generative capabilities from DMs.
+However, existing MU evaluation systems present several key challenges that can result in incomplete and inaccurate assessments.
+To address these issues, we propose UnlearnCanvas, a comprehensive high-resolution stylized image dataset that facilitates the evaluation of the unlearning of artistic styles and associated objects.
+This dataset enables the establishment of a standardized, automated evaluation framework with 7 quantitative metrics assessing various aspects of the unlearning performance for DMs.
+Through extensive experiments, we benchmark 9 state-of-the-art MU methods for DMs, revealing novel insights into their strengths, weaknesses, and underlying mechanisms.
+Additionally, we explore challenging unlearning scenarios for DMs to evaluate worst-case performance against adversarial prompts, the unlearning of finer-scale concepts, and sequential unlearning.
+We hope that this study can pave the way for developing more effective, accurate, and robust DM unlearning methods, ensuring safer and more ethical applications of DMs in the future.
+The dataset, benchmark, and codes are publicly available at https://unlearn-canvas.netlify.app/.
 1 Introduction
-In recent years, the field of text-to-image technology, epitomized by Diffusion Models (DMs), has seen remarkable progress, demonstrating an unprecedented ability to generate high-resolution, high-quality images across a wide array of subjects (Rombach et al., 2022; Midjourney et al.,, 2023).
-These models, widely accessible to public, find applications in diverse sectors, including advertising, the creative arts (Fotor et al.,, 2022), and forensic sketching (LabLab AI et al.,, 2023).
-Central to the training of DMs is their reliance on the data source from the Internet (Schuhmann et al., 2022) to cover a wide range of content.
-However, this inclusiveness comes with challenges, including potential concerns about harmful generation (Schramowski et al., 2023), copyright issues (Krietzberg, 2023), and biases/stereotypes (Vinker et al., 2023; Turk, 2023).
-Figure 1: An illustration of the task MU for DMs.
-The pretrained model contains the generation capability of different concepts in different domains.
-MU aims to erase the generation of a certain unlearning target concept without harming the generation of others.
-To alleviate the negative social impacts associated with DMs, machine unlearning (MU) techniques are catching increasing attention and have been studied in the field of vision generative models (Schramowski et al., 2023; Gandikota et al., 2023a, b; Kumari et al., 2023; Fan et al., 2023; Heng & Soh, 2023).
-MU for DMs (see Fig.
-1 for an illustration) aims to erase the model’s capacity to generate images related to a specific concept, essentially erasing that concept from the pre-trained model.
-This concept can be provided in the form of text prompts (i.e., text conditions of DMs), covering a wide range from concrete object entities to abstract notions, such as artistic styles.
-We refer readers to Sec.
-2 for more related work on MU for DMs.
-Unlike MU studies on image classification (Izzo et al., 2021; Golatkar et al., 2020; Becker & Liebig, 2022; Thudi et al., 2022; Warnecke et al., 2021; Jia et al., 2023), we lack evaluation benchmarks for MU in DMs.
-The current primary choice for assessing the performance of unlearned DMs (i.e., DMs post unlearning) is the Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023).
-This dataset consists of real-world image-to-text prompts designed for erasing concepts like ‘nudity’ and ‘violence’.
-However, I2P mainly contains isolated unlearning targets and focuses on safety issues related to image generation.
-To enhance the assessment of MU in DMs and establish a standardized evaluation framework, we propose to develop a new benchmark dataset, referred to as UnlearnCanvas, which is designed to evaluate the unlearning of artistic painting styles along with associated image objects.
-Unlike I2P, UnlearnCanvas can be used to assess the removal of copyrighted information, such as artistic style, and it allows for increased diversity by including various object classes and styles.
+The recent technological breakthroughs in text-to-image generation, driven by diffusion models (DMs), have shown an unprecedented capability to produce high-resolution, high-quality images across a diverse range of subjects [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
+These models have become widely accessible to the public and are applied across various sectors, including advertising, the creative arts [11], and forensic sketching [12].
+One reason for DMs being able to generate a broad spectrum of content is their reliance on diverse internet-sourced data [13].
+However, this inclusiveness comes with risks, such as harmful generation [14], copyright issues [15], and biases or stereotypes [16, 17].
+Refer to caption    Refer to caption
+(a) (b)
+Figure 1: (a) An illustration of MU for DMs.
+(b) Overview of experiment settings and benchmark results.
+This benchmark focuses on three categories of quantitative metrics: the unlearning effectiveness (UA, Rob., FU, SU); the retainability of innocent knowledge (IRA, CRA, FR, SR); and the image generation quality (FID).
+Results are normalized to $0\%\sim 100\%$ per metric.
+No single method excels across all metrics.
+See a summary of these metrics in Tab.
+A1 and more results in Sec.
+4.
+To alleviate the negative social impacts associated with DMs, machine unlearning (MU) techniques are catching increasing attention and have been studied in the field of text-to-image generation via DMs [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
+MU for DMs, also referred to as DM unlearning [27], aims to prevent the model from generating images when conditioned on an undesired concept (typically specified by a text prompt to be forgotten).
+Therefore, the problem of DM unlearning is also known as concept erasing [18, 19, 20, 21, 22, 23, 24, 25, 26].
+Based on the types of unlearning requests, the targeted concept to be erased can be diverse, including not-safe-for-work (NSFW) prompts [14, 23], concrete object entities [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], and copyrighted information like artistic styles [23, 24, 25, 26, 28, 29, 30, 32].
+In Fig.
+1 (a), we provide an illustration of DM unlearning by demonstrating the removal of an undesired artistic style from DM-generated images.
+Additionally, we refer readers to Sec.
+2 for more related work on DM unlearning.
+Figure 2: An illustration of machine unlearning using UnlearnCanvas.
+Concepts in the knowledge bank are categorized into different domains (style and object) and serve as potential unlearning targets.
+When one concept is unlearned, the rest concepts in both the same and different domains are required to be retained.
+Compared to MU for DMs, the unlearning studies in computer vision have primarily focused on image classification models [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
+Given an image classification dataset, benchmarking unlearning performance for discriminative models is simpler, because MU typically involves either class-wise forgetting [49, 35] or data-wise forgetting [27, 40].
+The former aims to eliminate the influence of a specific image class, while the latter removes the influence of specific data points from the entire training set.
+In contrast, DM unlearning is often applied to scenarios involving the removal of higher-level and more abstract concepts, rather than specific training data points, which can also be difficult to localize within the DM training set.
+The primary method for assessing the performance of DM unlearning is using the I2P (inappropriate image prompts) dataset [14].
+This dataset mainly focuses on the safety assessment of unlearned DMs (i.e., DMs post-unlearning), designed to erase specific harmful concepts like ‘nudity’ and ‘violence’.
+Although I2P specifies which concepts to unlearn, evaluating the unlearning effectiveness and generation retainability of unlearned DMs under I2P remains challenging.
+This difficulty arises primarily from the lack of ground-truth data or objective criteria to precisely define safety versus non-safety.
+Furthermore, I2P cannot evaluate the performance of DM unlearning in other scenarios, such as object unlearning [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and style unlearning [23, 24, 25, 26, 27, 28, 29, 30].
+To enhance the assessment of machine unlearning in DMs and establish a standardized evaluation framework, we propose the development of a new benchmark dataset, referred to as UnlearnCanvas.
+Unlike I2P, UnlearnCanvas is designed to evaluate the unlearning of artistic painting styles along with associated image objects (see an illustration in Fig.
+2).
+It also provides ground-truth data annotations to precisely define the criteria for the effectiveness of unlearning and the retained model utility post-unlearning (see a result overview in Fig.
+1 (b)).
 We summarize our contributions below.
-$\bullet$ We conduct a systematic review of existing MU methods for DMs and identified three unresolved challenges in their evaluations, such as limited quantitative assessments and a lack of retainability evaluation of unlearned DMs.
-$\bullet$ We propose UnlearnCanvas, a large-scale high-resolution stylized image dataset with 24,000 labeled images, 60+ styles, and 20+ object topics.
-Its style-object dual supervision and high stylistic consistency help us to improve the precision of MU evaluation for DMs.
-$\bullet$ We develop a standardized evaluation pipeline for MU in DMs under UnlearnCanvas, which measures unlearning effectiveness, generation retainability, generation quality, and unlearning efficiency.
-$\bullet$ We benchmark state-of-the-art MU methods for DMs using UnlearnCanvas, validating the significance of our proposed evaluation metrics and providing insights into the strengths, weaknesses, and unlearning mechanisms.
-2 Related Work and Preliminaries
-Machine unlearning (MU) for generative models in vision.
-MU was originally developed to mitigate the (potentially detrimental) influence of specific data points in a pretrained ML model, without necessitating a complete retraining of the model after removing these unlearning data (Cao & Yang, 2015; Bourtoule et al., 2021; Xu et al., 2023).
-The significance of MU has emerged with the purpose of data privacy protection, in response to regulations such as ‘the right to be forgotten’ (Hoofnagle et al., 2019).
+$\bullet$ We conduct a systematic review of existing MU methods for DMs and identified three unresolved challenges in their evaluations.
+We further provide an in-depth analysis of why a standardized benchmark for DM unlearning evaluation is crucial.
+$\bullet$ We propose UnlearnCanvas, a large-scale, high-resolution stylized image dataset with diverse styles and objects, to address current challenges in DM unlearning evaluation.
+Its dual supervision of styles and objects ensures stylistic consistency, enabling a comprehensive, standardized evaluation approach to precisely characterize unlearning efficacy, generation utility preservation, and efficiency.
+$\bullet$ We benchmark 9 state-of-the-art DM unlearning methods using UnlearnCanvas, covering standard assessments and examining more challenging scenarios, such as adversarial robustness, grouped object-style unlearning, and sequential unlearning.
+This comprehensive evaluation provides previously unknown insights into their strengths and weaknesses.
+2 Related Work, Problem Statement, and Open Challenges
+Related work.
+MU (machine unlearning) was originally developed to mitigate the (potentially detrimental) influence of specific data points in a pretrained ML model, without necessitating a complete retraining of the model after removing these unlearning data [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
+The significance of MU has emerged with the purpose of data privacy protection, in response to regulations such as ‘the right to be forgotten’ [53].
 However, it has rapidly gained recognition for its role in promoting security, safety, and trustworthiness of ML models.
-Examples include defense against backdoor attacks (Liu et al., 2022; Jia et al., 2023), fairness enhancement (Chen et al., 2023; Oesterling et al., 2023), and controllable federated learning (Halimi et al., 2022; Jin et al., 2023).
-More recently, unlearning has proven crucial in text-to-image generation via DMs to prevent the production of harmful, private, or illegal content (Gandikota et al., 2023a; Zhang et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023; Kumari et al., 2023; Heng & Soh, 2023).
-For example, DMs have demonstrated a susceptibility to generating NSFW (Not Safe For Work) images when conditioned on inappropriate text prompts (e.g., ‘nudity’ and ‘violence’) (Schramowski et al., 2023; Zhang et al., 2023b; Chin et al., 2023).
-Similarly, MU has also found applications in preventing the generation of copyrighted information, such as the misuse of artistic painting styles (Gandikota et al., 2023a; Zhang et al., 2023a).
-In summary, MU can be regarded as a (post-training) model edit operation, aimed at preventing DMs from producing undesired or inappropriate images.
-Our setup and objective.
+Examples include defense against backdoor attacks [38, 54], fairness enhancement [55, 56], and controllable federated learning [57, 58].
+More recently, DM unlearning has proven crucial in text-to-image generation to prevent the production of harmful, private, or illegal content [23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
+For example, DMs have demonstrated a susceptibility to generating NSFW (not-safe-for-work) images when conditioned on inappropriate text prompts (e.g., ‘nudity’ and ‘violence’) [14, 59, 60].
+Similarly, MU has also found applications in preventing the generation of copyrighted information, such as the misuse of artistic painting styles [23, 28].
+In brief, DM unlearning can be regarded as a model edit operation, aimed at preventing DMs from producing undesired or inappropriate images.
+As unlearning approaches continue to evolve, the need for benchmarking their performance becomes increasingly critical.
+Standard datasets like CIFAR-10 [61] and ImageNet [62] have long been used to evaluate unlearning performance when forgetting a specific subset of data points [27, 40] or a targeted image class [49, 35].
+Additionally, a recent benchmark specialized for unlearning in face recognition has been developed [63].
+However, crafting effective benchmarks for generative models introduces additional challenges.
+In the language domain, recent initiatives have specifically aimed to evaluate the efficacy of MU for large language models (LLMs).
+Notable examples include TOFU [64], which explores the unlearning of fictitious elements, and WMDP [65], aimed at removing hazardous knowledge from LLMs.
+In contrast, to the best of our knowledge, there is currently no effective benchmark dataset designed for DM unlearning in text-to-image generation.
+The predominant dataset, I2P [14], targets the removal of harmful concepts for safe image generation.
+However, I2P only covers a specialized unlearning scenario and lacks precise evaluation criteria to support a comprehensive, quantitative, and precise assessment of unlearning performance.
+This gap inspires us to develop a new benchmark dataset for DM unlearning.
+Table 1: Overview of existing MU evaluations for DMs, which covers two unlearning scenarios: object unlearning and style unlearning.
+Each case is demonstrated by the number of targeted unlearning concepts (target #), qualitative evaluation (Qual.) via generated image visualization, and quantitative evaluation (Quant.) of unlearning effectiveness (UE) and retainability (RT).
+Related Work Object Unlearning Style Unlearning
+Target # Quant. Qual. Target # Quant. Qual.
+UE RT UE RT
+ESD [23] 11 5
+CA [16] 4 4
+FMN [28] 7 1
+UCE [24] 11 7
+SA [29] 31 1
+SalUn [27] 11 0
+SPM [26] 24 5
+SEOT [30] 22 2
+EDiff [31] 2 0
+SHS [32] 1 0
+Problem statement.
 Let us denote a pretrained DM by $\bm{\theta}_{\mathrm{o}}$, capable of synthesizing high-quality images conditioned on a provided text prompt $c$ (e.g., ‘A painting of a cat in Van Gogh style’).
-In this work, we focus on the task of generating images in various painting styles.
-When there is a request to adapt the model $\bm{\theta}_{\mathrm{o}}$ to prevent the generation of images, e.g., in the ‘Van Gogh’ style, the MU problem arises, with the ‘Van Gogh’ style being the specified unlearning target, also known as the erasing concept (Gandikota et al., 2023a).
+When there is a request to adapt the model $\bm{\theta}_{\mathrm{o}}$ to prevent generating images in a specific concept, such as the ‘Van Gogh’ painting style, the MU problem arises.
+Here, the ‘Van Gogh’ style becomes the specified unlearning target, also known as the erasing concept [23].
 See Fig.
-1 for a visual representation of the concept.
-In formulating MU methods, the traditional setup presupposes access to the training dataset $\mathcal{D}$, where $\mathcal{D}_{\mathrm{f}}\subset\mathcal{D}$ represents the forget set that revolves around the unlearning target.
-A commonly used gold standard of MU is to retrain the model from scratch over the retain set $\mathcal{D}_{\mathrm{r}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$.
-However, commencing the training process from scratch becomes practically unfeasible because of (1) the substantial computational expense it entails, (2) the difficulty in accurately identifying $\mathcal{D}_{\mathrm{f}}$, and (3) the absence of access to the original training data source.
-To overcome the aforementioned difficulties (1)-(3), the current research mainstream (Gandikota et al., 2023a; Kumari et al., 2023; Zhang et al., 2023a; Gandikota et al., 2023b; Heng & Soh, 2023; Fan et al., 2023) has focused on developing approximate yet efficient unlearning methods as practical alternatives to full retraining.
-The approximation introduced by current unlearning methods raises the question that we aim to address in this work:
-How can we accurately assess unlearning performance and establish a benchmark for existing unlearning methods?
-3 Unresolved Challenges in MU Evaluation
-This work will develop a new benchmark dataset, referred to as UnlearnCanvas (detailed in Sec.
-4), focusing on the task of unlearning artistic painting styles in conjunction with associated image objects.
-With the use of UnlearnCanvas, a text prompt for DMs can be constructed as ‘A painting of a {object} in {style}’.
-And the unlearning target can be explicitly designated as either ‘style’ or ‘object’.
-To be specific, UnlearnCanvas is targeted for the open challenges (C1-C3) when building a rigorous quantitative evaluation framework for MU in DMs.
-These challenges are consolidated through a careful examination of existing studies for object and/or style unlearning in Tab.
+1-(a) for a visual representation of this unlearning problem.
+In this work, we specify the unlearning target as artistic styles and/or object classes to encompass the settings of style unlearning and object unlearning as described in the literature [23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
+Motivation: Unresolved challenges in MU evaluation.
+When building a rigorous quantitative evaluation framework for DM unlearning, we identified three urgent challenges (C1)-(C3) after carefully examining existing studies as shown in Tab.
 1.
-Table 1: An overview of the scope and evaluation methods of the existing MU methods for DMs. The symbol denotes neither objects nor styles are covered.
-Related work Unlearning targets Quantitative evaluation
-Unlearning
-effectiveness
-Retainability
-ESD (Gandikota et al., 2023a)
-5 artistic styles
-11 objects
-objects only objects only
-CA (Kumari et al., 2023)
-4 artistic styles
-4 objects
-objects only
-FMN (Zhang et al., 2023a)
-1 artistic style
-7 objects
-objects only
-UCE (Gandikota et al., 2023b)
-7 artistic settings [^1] UCE adopted a multi-concept unlearning setting, where different numbers of styles are unlearned at the same time and total up to 1000 styles are considered.
-11 objects
-objects only objects only
-SA (Heng & Soh, 2023)
-1 artistic style
-31 objects
-objects only
-SalUn (Fan et al., 2023)
-No artistic styles
-11 objects
-objects only objects only
-(C1) The absence of a diverse yet cohesive unlearning target repository.
+(C1) The absence of a consensus on a diverse unlearning target test repository.
 As shown in Tab.
-1, the assessment of MU for DMs, in terms of both effectiveness (concept erasure performance) and retainability (preserved generation quality under non-forgotten concepts), is typically conducted using manually selected targets, often chosen from a limited pool of unlearning targets.
-A critical aspect missing from MU evaluation is the recognition that unlearning targets related to styles and objects should not be isolated within the dataset used for MU assessment.
-Instead, they should be considered in a manner that exposes their potential interconnected influences on other targets.
-Table 2: Classification accuracy of a style classifier (ViT-L/16) finetuned on WikiArt (Phillips & Mackintosh, 2011).
-The classifier is tested on the test set and the image set generated by SD v1.5 and v2.0 with the prompt ‘An painting in {artist} style’, where the {artist} is traversed over the WikiArt labels (10 images/prompt).
-WikiArt
-Train Set
-WikiArt
-Test Set
-Images by
-SD v1.5
-Images by
-SD v2.0
-Accuracy 99.8% 85.4% 56.7% 41.2%
-(C2) The lack of precision in evaluation.
-Artistic styles can be challenging to precisely define and distinguish, which makes quantifying their impact on unlearning effectiveness and retainability difficult.
-To illustrate this challenge, we conducted a preliminary investigation in Tab.
-2 using a ViT-based classifier trained on the WikiArt dataset (Phillips & Mackintosh, 2011).
-This dataset categorizes paintings into 139 distinct styles, each annotated with the creator’s name.
-Despite training, the ViT-large model achieved only an $84.32\%$ accuracy on the test dataset.
-This result underscores the inherent complexity of classifying artistic styles, causing inaccuracies in quantitative MU evaluations.
-The classifier’s suboptimal accuracy on DMs generated images further increases the prediction error (around $50\%$ accuracy).
-This complicates precise quantitative evaluations.
-(C3) The lack of a systematic study on ‘retainability’ of DMs post unlearning.
-As indicated by the ‘retainability’ column in Tab.
-1, assessing the capacity of unlearned DMs to maintain image generation under innocent concepts is notably lacking.
-However, retainability is a critical indicator of potential ‘side effects’ of an unlearning method.
-In conjunction with challenge (C1), the inclusion of diverse unlearning targets, encompassing both styles and objects, allows for the evaluation of retainability from various perspectives.
-For instance, when unlearning an artistic style, it is crucial to consider the retainability of image generation abilities for other styles as well as for objects.
-The challenges (C1)-(C3) outlined above have driven us to develop UnlearnCanvas as a solution for the systematic and comprehensive evaluation of MU in DMs.
-We will show in Sec.
-4 that UnlearnCanvas includes a diverse range of artistic styles and objects.
-Within this dataset, we can explore the interrelationships between paired image styles and image objects to address (C1).
-Each style category in UnlearnCanvas maintains high intra-style consistency, facilitating accurate style and object classification and enabling post-unlearning quantitative evaluation for (C2).
-To address (C3), we can evaluate two types of retainability, providing a precise and comprehensive assessment of the potential side effects induced by MU methods.
+1, assessments of DM unlearning in various studies, in terms of both effectiveness (i.e., concept erasure performance) and retainability (i.e., preserved generation quality under non-forgotten concepts), typically use manually-selected unlearning targets from a limited pool and focus on object-centric evaluation.
+Precise evaluation for style unlearning and grouped object-style unlearning is lacking.
+Although the existing dataset WikiArt [66] features a collection of real-world artworks by various artists, it does not apply to object unlearning.
+(C2) The lack of a systematic study on ‘retainability’ of DMs post-unlearning.
+As highlighted in the ‘retainability’ column of Tab.
+1, there is a notable deficiency in the quantitative assessment of retainability for DMs post-unlearning.
+Retainability is crucial for measuring the potential side effects of a MU method.
+For instance, when unlearning an artistic style, it is essential to assess the DM’s ability to generate images in other styles and all objects.
+This makes the WikiArt dataset less suitable for benchmarking DM unlearning, as it lacks object labels needed to characterize the side effects of style unlearning on object recognition.
+(C3) The precision challenge in evaluating DM-generated images.
+Artistic styles are inherently complex to define and differentiate precisely, complicating the quantitative evaluation of forgetting and retaining performance in style unlearning.
+For example, the WikiArt dataset was recently used in [67] to train a classifier for detecting copyright infringement, merely achieving an accuracy of 72.80%.
+The difficulty in precisely recognizing artistic styles in WikiArt may further hamper the evaluation of DMs’ post-unlearning generation.
+Tab.
+A3 illustrates the challenge of accurately recognizing the styles of images generated by DMs, even when finetuned on WikiArt.
+This is evidenced by the test-time classification accuracy using the style classifier, ViT-L/16 [68], finetuned on WikiArt.
+As we can see, only about half of the images are correctly classified into their corresponding style classes, e.g., 56.7% accuracy for classification on generated images using finetuned stable diffusion (SD) v2.0 [1], which is significantly lower than that on test images from the original WikiArt (85.4%).
+The above challenges (C1)-(C3) drive us to develop the UnlearnCanvas dataset as a solution for the systematic and comprehensive evaluation of DM unlearning in Sec. 3.
 3 Our Proposal: UnlearnCanvas Dataset
@@
 3 for a schematic overview.
+More information about the dataset is provided in Appx.
+A.
+Figure 3: Illustration of curating UnlearnCanvas.
 Seed image collection.
 UnlearnCanvas includes images in $60$ unique artistic styles across $20$ distinct objects.
-The style categories are derived from a set of high-resolution photos, termed seed images.
-These seed images consist of original real-world photos from Pexels (Pexels et al.,, 2023).
-In total, there are 400 seed images across 20 object classes.
+The style categories are built upon a set of high-resolution seed images, given by real-world photos from Pexels [69].
+There are $20$ seed images collected for each of the $20$ object classes.
+See step 1 in Fig.
+3.
 Image stylization.
-After collecting the seed images, the next step in dataset development is the stylization process.
-During this phase, each seed image is transformed into 60 predetermined artistic styles.
-The goal is to ensure that the resulting images maintain high stylistic consistency while preserving the content details present in the seed images.
-This stylization process is carried out using the services provided by (Fotor et al.,, 2022).
-Once all the seed images have been stylized in all 60 styles, the dataset is constructed with super-high-resolution images and organized in a hierarchical structure that balances both style and object categories.
-Figure 2: An illustration of the key steps when curating UnlearnCanvas and its key features.
-Advantages of UnlearnCanvas. Next, we will highlight the advantages (A1-A3) of UnlearnCanvas to address the MU evaluation challenges (C1-C3) discussed in Sec. 3.
+After collecting the seed images, the next step is the stylization process.
+During this phase, each seed image is transformed into 60 predetermined artistic styles, ensuring high stylistic consistency while preserving the original content details.
+This stylization process is carried out using services provided by Fotor [11].
+Once all seed images have been stylized in all 60 styles, the dataset is constructed with super-high-resolution images and organized in a hierarchical structure that balances both style and object categories.
+See step 2 in Fig.
+3.
+Advantages of UnlearnCanvas. Next, we will elaborate on the advantages (A1)-(A3) of UnlearnCanvas to address the MU evaluation challenges (C1)-(C3) discussed in Sec. 2.
 (A1) Style-object dual supervision enables a rich unlearning target bank for comprehensive evaluation.
-UnlearnCanvas offers a diverse range of unlearning targets by annotating each image with both style and object labels.
-This is in response to (C1).
-We have categorized the unlearning targets into two main domains: a style domain comprising $60$ artistic styles, and an object domain including $20$ object classes.
-This setup allows us to investigate the impact of style unlearning on the fidelity of generated objects in images and vice versa.
-Besides its role in MU evaluation, UnlearnCanvas could also serve as an annotated testbed for assessing style transfer accuracy and visual in-context learning performance; see details in Appx.
-C.
+To address (C1), UnlearnCanvas offers a diverse range of unlearning targets by annotating each image with both style and object labels.
+We categorized the unlearning targets into two main domains: a style domain comprising 60 styles, and an object domain with 20 object classes.
+This setup allows us to investigate the impact of style unlearning on the fidelity of object generation and vice versa.
+(A2) Enabling both ‘in-domain’ and ‘cross-domain’ retainability analyses for MU evaluation.
+To address (C2), UnlearnCanvas facilitates us to assess DMs’ retainability post-unlearning from two perspectives: in-domain retainability, which evaluates the DM’s generation performance within the same domain as the unlearning target, and cross-domain retainability, which assesses the model’s generation ability under conditions from a different domain.
+For example, as depicted in Fig.
+4, we can define the ‘domain’ as either ‘artistic style’ or ‘object class’.
+If the unlearning target is the ‘Van Gogh’ style, then generating images in all styles except ‘Van Gogh’ within the same object class (e.g., ‘Dog’ in Fig.
+4) represents the in-domain case.
+Conversely, cross-domain retainability refers to generating images of different object classes (e.g., ‘Horses’ in Fig.
+4), corresponding to an object domain different from the unlearning target within the style domain.
+Figure 4: Illustration of in-domain and cross-domain retainability evaluation, with the Van Gogh style as the unlearning target. ✓ and ✗ indicate satisfactory and undesired results post unlearning.
 (A3) High stylistic consistency ensures precise style definitions and enables accurate quantitative evaluations.
-In UnlearnCanvas, images labeled with the same style exhibit high intra-style consistency and distinct inter-style differences, as exemplified in Fig.
-2.
+In UnlearnCanvas, images labeled with the same style exhibit high intra-style consistency and distinct inter-style differences.
 This facilitates the distinction between different styles, effectively addressing challenge (C2).
 In a similar study to that presented in Tab.
-2, we observe that the styles within UnlearnCanvas can be readily classified using ViT-Large (Dosovitskiy et al., 2020), achieving an accuracy of $98\%$ for images generated by SD finetuned on UnlearnCanvas.
-We refer readers to Tab.
-A2 for detailed results.
-Figure 3: Evaluation metric illustration, with unlearning the ‘Van Gogh style’ as a case study. ✓ and ✗ denote desired and undesired results post unlearning.
-(A3) Enabling in-depth retainability analyses for MU evaluation.
-Using UnlearnCanvas, we can assess DMs’ retainability post-unlearning from two perspectives: in-domain retainability, which evaluates the model’s performance on test-time generations within the same domain as the unlearning target, and cross-domain retainability, which assesses the model’s ability to generate concepts from a different domain.
-For example, as depicted in Fig.
-3, we can define the ‘domain’ as either ‘artistic style’ or ‘object class’.
-If the unlearning target is the ‘Van Gogh’ style (i.e., the concept to be erased), then generating images in all styles except ‘Van Gogh’ but within the same object class (i.e., ‘Dog’ in Fig.
-3) represents the in-domain case.
-Conversely, out-of-domain refers to generating images with different object classes (e.g., ‘Horses’ in Fig.
-3), i.e., objects different from the one associated with the unlearning target.
-Figure 4: Image examples with the label “Van Gogh Style” from WikiArt (Phillips & Mackintosh, 2011) and UnlearnCanvas.
-Images of the same artistic style in UnlearnCanvas exhibit high stylistic consistency compared to WikiArt.
-Figure 5: An illustration of the evaluation pipeline proposed in this work using UnlearnCanvas when unlearning a specific target concept ‘Van Gogh Style’.
-Unlearning performances (including the unlearning effectiveness and retainability) are quantitatively assessed (marked in blue) to accurately reflect the unlearning performance portrait of each method.
-The unlearning target of the pipeline could traverse all the styles and objects to achieve a complete and comprehensive evaluation.
-UnlearnCanvas vs. WikiArt.
-To the best of our knowledge, WikiArt (Nichol, 2016) is the most relevant baseline dataset to ours.
-In Tab.
-A1, we provide a direct comparison of the key attributes of these two datasets.
-UnlearnCanvas differs from WikiArt in the following aspects.
-First, UnlearnCanvas includes a greater number of high-resolution images ($15$M) compared to WikiArt ($2$M), a factor that may enhance the training of state-of-the-art DMs.
-Second, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness, as illustrated in Fig.
-4, where images labeled with ‘Van Gogh Style’ from both datasets are compared.
-In UnlearnCanvas, the images exhibit high stylistic consistency, while WikiArt lacks the necessary clarity for precise assessment.
-This will hamper the MU evaluation as discussed in the challenges (C2) and (C3).
-5 Evaluation Pipeline
-In this section, we introduce the evaluation pipeline and the benchmarked MU methods for UnlearnCanvas.
+A3, we observe that the styles within UnlearnCanvas can be readily classified using ViT-Large [68], achieving an accuracy of 98% for images generated by DM finetuned on UnlearnCanvas.
+For detailed results, please refer to Tab.
+A4.
+The above advantages (A1)-(A3) make UnlearnCanvas uniquely suitable for assessing the performance of DM unlearning.
+For instance, UnlearnCanvas includes a greater number of high-resolution images (15M) compared to the existing dataset WikiArt (2M).
+Moreover, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness.
+See Appx.
+C for more detailed comparisons.
+Evaluation pipeline via UnlearnCanvas.
+We next introduce how UnlearnCanvas can be used to benchmark the performance of DM unlearning and the corresponding evaluation metrics.
 Fig.
-5 illustrates the proposed evaluation pipeline for MU in DMs with a designated unlearning target.
+5 illustrates the proposed evaluation pipeline with a designated unlearning target.
 It comprises four phases (I-IV) to evaluate unlearning effectiveness, retainability, generation quality, and efficiency.
 Phase I: Testbed preparation (Fig.
-5a).
-Prior to unlearning, we commence by fine-tuning a specific DM, given by SD (StableDiffusion) v1.5, on UnlearnCanvas for text-to-image generation.
-We also train a vision transformer (ViT-Large) (Dosovitskiy et al., 2020) for the purposes of style and object classification after unlearning.
+5 (a)).
+Prior to unlearning, we commence by finetuning a specific DM, given by SD (StableDiffusion) v1.5, on UnlearnCanvas for text-to-image generation.
+We also train a vision transformer (ViT-Large) [68] for the purposes of style and object classification after unlearning.
 We ensure that both the SD model and classifiers attain high performance in image generation and classification across all styles and objects.
 See detailed testbed preparation in Appx.
-A.
+B.
 Phase II: Machine unlearning (Fig.
-5b).
-Next, we utilize the selected MU methods for benchmarking to update the DM acquired in Phase I, aiming to eliminate a designated unlearning target (e.g., ‘Van Gogh style’).
-This results in the unlearned model (denoted as ${\bm{\theta}}_{\mathrm{u}}$), which is the DM updated to avoid generating the unlearning target.
-This unlearned DM will be used for unlearning performance evaluation.
+5 (b)).
+Next, we utilize a given MU method to update the DM acquired in Phase I, aiming to eliminate a designated unlearning target (e.g., Van Gogh style).
+This results in the unlearned DM (denoted as ${\bm{\theta}}_{\mathrm{u}}$), which avoids generating images in response to Van Gogh style-specific prompts.
 Phase III: Answer set generation (Fig.
-5c).
-Further, we utilize the unlearned model ${\bm{\theta}}_{\mathrm{u}}$ to generate a set of images conditioned on both the unlearning-related prompts and other innocent prompts.
-To comprehensively evaluate unlearning performance, three types of answer sets are generated:
-Unlearning effectiveness: Images are generated in response to prompts that involve the unlearning target (e.g., any prompt like ‘An image in Van Gogh style’ for the case in Fig.
+5 (c)).
+Further, we utilize the unlearned model ${\bm{\theta}}_{\mathrm{u}}$ to generate a set of images conditioned on both the unlearning-related prompts and other benign prompts.
+To comprehensively evaluate unlearning performance, three types of answer sets are generated.
+(a) To assess unlearning effectiveness, images are generated in response to prompts that involve the unlearning target (e.g., any prompt like ‘An image in Van Gogh style’ for the case in Fig.
 5).
-A successfully unlearned model will avoid generating images that match the unlearning target.
-In-domain retainability: Images are generated with prompts in the same ‘domain’ as the unlearning target (e.g., style-related prompts for style unlearning), where the prompts can cover all the other styles provided in UnlearnCanvas except the one to be unlearned, e.g., ‘Van Gogh style’ in Fig.
-5.
-This assesses the unlearned model’s ability to retain generation quality within the same domain.
-Cross-domain retainability: Images are generated with prompts in other domains.
+(b) To assess in-domain retainability (i.e., the unlearned model’s ability to retain generation quality within the same domain), images are generated with prompts in the same domain as the unlearning target (e.g., style-related prompts for style unlearning).
+These prompts cover all other styles provided in UnlearnCanvas except the one to be unlearned.
+(c) To measure cross-domain retainability, images are generated with prompts in other domains.
 For instance, in the case of style unlearning as shown in Fig.
@@
 Prompts related to objects are used to evaluate the DM’s cross-domain generation capability after style unlearning.
-Table 3: Performance overview of different MU methods evaluated with UnlearnCanvas dataset.
-The performance assessment includes unlearning accuracy (UA), in-domain retain accuracy (IRA), cross-domain retain accuracy (CRA), and FID.
-Symbols $\uparrow$ or $\downarrow$ denote whether larger or smaller values represent better performance.
+Phase IV: MU performance assessment (Fig.
+5 (d)).
+After Phase III, the answer set undergoes style/object classification for unlearning performance assessment.
+This classification results in three quantitative metrics ①-③.
+① Unlearning accuracy (UA): This represents the proportion of images generated by the unlearned DM using the unlearning target-related prompt that is not correctly classified into the corresponding class.
+A higher UA indicates better unlearning performance in preventing image generation from the unlearning target-related prompts.
+② In-domain retain accuracy (IRA): This is the classification accuracy of images generated by the unlearned DM using innocent prompts (not relevant to the unlearning target) within the same domain.
+③ Cross-domain retain accuracy (CRA): Similar to IRA, this is the classification accuracy of images generated by the unlearned DM using innocent prompts in different domains.
+In addition to the accuracy metrics, we also use the FID score (④) to evaluate the distribution-wise generation quality of the unlearned DM.
+Furthermore, we monitor the efficiency of the unlearning process, considering factors such as run-time (⑤), storage space requirements (⑥), and memory costs (⑦).
+Figure 5: An illustration of the evaluation pipeline proposed in this work using UnlearnCanvas when unlearning a specific target concept ‘Van Gogh Style’.
+Unlearning performances (including the unlearning effectiveness and retainability) are quantitatively assessed (marked in blue) to accurately reflect the unlearning performance portrait.
+The unlearning target of the pipeline could traverse all the styles and objects to achieve a comprehensive evaluation.
+DM unlearning methods to be benchmarked.
+In this work, we assess 9 most recently proposed MU methods for DMs, including ESD [23], FMN [28], UCE [24], CA [25], SalUn [27], SEOT [30], SPM [26], EDiff [31], and SHS [32].
+We remark that SA [29], a recently proposed method, is excluded from our evaluation due to its excessive computational resource requirements and time consumption.
+Unless specified otherwise, SD v1.5 is the model used for unlearning.
+Detailed training settings for each method can be found in Appx.
+B.
+4 Experiment Results
+In this section, our benchmarking results are divided into two main parts.
+In Sec.
+4.1, we comprehensively evaluate the existing DM unlearning methods on the established tasks of style and object unlearning.
+Extensive studies following the proposed evaluation pipeline (Fig.
+5) reveal that focusing solely on unlearning effectiveness can lead to a biased perspective for DM unlearning if retainability is not concurrently assessed (Tab.
+2 & Fig.
+A2).
+We also show that preserving CRA (cross-domain retainability) is more challenging than IRA (in-domain retainability), highlighting a gap in the current literature (Fig.
+6).
+Furthermore, we find that different DM unlearning methods exhibit distinct unlearning mechanisms by examining their unlearning directions (Fig.
+7).
+In Sec.
+4.2, we introduce more challenging unlearning scenarios.
+First, we assess the robustness of current DM unlearning methods against adversarial prompts crafted based on (Fig.
+8).
+Additionally, we examine the performance when facing unlearning targets with finer granularity, formed by style-object concept combinations (Tab.
+3).
+Further, we leverage UnlearnCanvas to provide insights into DM unlearning in a sequential unlearning fashion (Tab.
+A5).
+4.1 Benchmarking Current DM Unlearning Methods for Style and Object Unlearning
+Overall performance of DM unlearning.
+In Tab. 2, we provide an overview of the performance of existing unlearning methods, using the evaluation metrics associated with UnlearnCanvas.
+Table 2: Performance overview of different DM unlearning methods evaluated on UnlearnCanvas.
+The performance metrics include UA (unlearning accuracy), IRA, CRA, and FID.
+The symbols $\uparrow$ or $\downarrow$ indicate whether a higher or lower value represents better performance.
 Results are averaged over all the style and object unlearning cases.
-The best performance regarding each metric is highlighted in bold.
+The best performance for each metric is highlighted in green, while significantly underperforming results are marked in red, indicating areas needing improvement for existing DM unlearning methods.
 Method Effectiveness Efficiency
@@
 UA ($\uparrow$) IRA ($\uparrow$) CRA ($\uparrow$) UA ($\uparrow$) IRA ($\uparrow$) CRA ($\uparrow$) (s) ($\downarrow$) (GB) ($\downarrow$) (GB) ($\downarrow$)
-ESD (Gandikota et al., 2023a) $\mathbf{98.58}\%$ $80.97\%$ $93.96\%$ $92.15\%$ $55.78\%$ $44.23\%$ $65.55$ 6163 17.8 4.3
-FMN (Zhang et al., 2023a) $88.48\%$ $56.77\%$ $46.60\%$ $45.64\%$ $90.63\%$ $73.46\%$ $131.37$ $\mathbf{350}$ 17.9 4.2
-UCE (Gandikota et al., 2023b) $98.40\%$ $60.22\%$ $47.71\%$ $\mathbf{94.31}\%$ $39.35\%$ $34.67\%$ $182.01$ 434 $\mathbf{5.1}$ $\mathbf{1.7}$
-CA (Kumari et al., 2023) $60.82\%$ $\mathbf{96.01}\%$ $92.70\%$ $46.67\%$ $90.11\%$ $81.97\%$ $\mathbf{54.21}$ 734 10.1 4.2
-SalUn (Fan et al., 2023) $86.26\%$ $90.39\%$ $\mathbf{95.08}\%$ $86.91\%$ $\mathbf{96.35}\%$ $\mathbf{99.59}\%$ $61.05$ 667 30.8 4.0
-Phase IV: MU performance assessment (Fig.
-5d).
-After Phase III, the answer set undergoes style/object classification for unlearning performance assessment.
-This classification results in three quantitative metrics ①-③.
-Unlearning accuracy (UA): It represents the proportion of images generated by the unlearned DM using the unlearning target-related prompt (e.g., the Van Gogh style prompt in Fig.
-5), which are not correctly classified into the corresponding class for the case.
-Therefore, a higher UA indicates better unlearning performance in preventing image generation from the unlearning target-related prompts.
-In-domain retain accuracy (IRA): It is given by the classification accuracy of images generated by the unlearned DM using innocent prompts (not relevant to the unlearning target) in the same domain, e.g., all styles other than the Van Gogh style in Fig.
-5.
-Cross-domain retain accuracy (CRA): Similar to IRA, it is given by the classification accuracy of images generated by the unlearned DM using innocent prompts but in different domains, e.g., all object categories in Fig.
-5.
-In addition to the above accuracy metrics, we also utilize the FID score (④) to evaluate the distribution-wise generation quality of the unlearned DM.
-Furthermore, we will also monitor the efficiency of the unlearning process, taking into account factors such as run-time (⑤), storage space requirements (⑥), and memory costs (⑦).
-In total, the performance evaluation involves 7 metrics (①-⑦).
-MU methods for DMs to be benchmarked.
-In this work, we assess the 5 most recently proposed MU methods for DMs, including ESD (Gandikota et al., 2023a), FMN (Zhang et al., 2023a), UCE (Gandikota et al., 2023b), CA (Kumari et al., 2023), and Salun (Fan et al., 2023).
-We remark that SA (Heng & Soh, 2023), a recently proposed method, is excluded from our evaluation due to its excessive computational resource requirements and time consumption.
-Detailed training settings for each unlearning method can be found in Appx.
-A.
-Figure 6: Performance visualization of Table 3.
-For UA, IRA, and CRA, the results are averaged over all the style and object unlearning scenarios.
-Other metrics undertake the inverse operation as a smaller values represent better performance.
-Results are normalized to $0\%\sim 100\%$ per metric.
-Figure 7: Left: Heatmap visualization of the unlearning and retain accuracy of ESD on UnlearnCanvas.
+ESD [23] ${98.58}\%$ $80.97\%$ $93.96\%$ $92.15\%$ ${55.78\%}$ ${44.23\%}$ $65.55$ 6163 17.8 4.3
+FMN [28] $88.48\%$ ${56.77\%}$ ${46.60\%}$ ${45.64\%}$ $90.63\%$ $73.46\%$ 131.37131.37{131.37}131.37 $350$ 17.9 4.2
+UCE [24] ${98.40}\%$ ${60.22\%}$ ${47.71\%}$ ${94.31}\%$ ${39.35\%}$ ${34.67\%}$ 182.01182.01{182.01}182.01 434 5.15.1{5.1}5.1 1.71.7{1.7}1.7
+CA [25] ${60.82\%}$ ${96.01}\%$ $92.70\%$ ${46.67\%}$ $90.11\%$ $81.97\%$ 54.2154.21{54.21}54.21 734 10.1 4.2
+SalUn [27] $86.26\%$ $90.39\%$ $95.08\%$ $86.91\%$ ${96.35}\%$ ${99.59}\%$ $61.05$ 667 30.8 4.0
+SEOT [30] ${56.90\%}$ $94.68\%$ $84.31\%$ ${23.25\%}$ $95.57\%$ $82.71\%$ 62.38 9595{95}95 7.34 0.0
+SPM [26] ${60.94\%}$ $92.39\%$ $84.33\%$ $71.25\%$ $90.79\%$ $81.65\%$ 59.79 29700 6.9 0.0
+EDiff [31] $92.42\%$ ${73.91\%}$ ${98.93}\%$ $86.67\%$ $94.03\%$ ${48.48\%}$ 81.42 1567 27.8 4.0
+SHS [32] $95.84\%$ $80.42\%$ ${43.27\%}$ $80.73\%$ $81.15\%$ ${67.99\%}$ 119.34 1223 31.2 4.0
+First, as seen in Tab.
+2, retainability is essential for a comprehensive assessment of DM unlearning.
+For example, in the scenario of style unlearning, ESD and UCE achieve similar UA (both around $98\%$), but their retainability (IRA and CRA) differs significantly, with one over $80\%$ and the other around $60\%$.
+Therefore, relying solely on UA can provide a skewed view of the performance of DM unlearning.
+Second, we observe that CRA is harder to retain than IRA.
+For example, UCE exhibits a significant gap between its IRA and CRA, both in style ($60.22\%$ vs.
+$47.71\%$) and object ($39.35\%$ vs.
+$34.67\%$) unlearning scenarios.
+Similar patterns are observed with other methods like FMN, CA, SEOT, SPM, and SHS.
+This pattern suggests that while MU methods are somewhat effective at preserving concepts within the same domain, they face greater challenges in maintaining performance on unlearning target-unrelated concepts across domains.
+This issue has been overlooked in previous studies due to the absence of a multi-label dataset like UnlearnCanvas for systematic unlearning analyses.
+Third, we find that a single unlearning method can perform differently across various domains, and no single method excels in all aspects.
+For example, FMN achieves a UA of 88.48% in style unlearning but a much lower UA of 45.64% in object unlearning.
+A similar phenomenon can be observed with the unlearning methods SEOT, SPM, and SHS, which exhibit significantly large performance gaps between style and object unlearning.
+Furthermore, each method exhibits unique strengths and weaknesses.
+For example, while ESD typically shows improvement in UA, it may lag in terms of IRA and CRA, particularly in the case of object unlearning.
+Some methods, such as SEOT and SPM, achieve extremely high storage efficiency, requiring no additional storage, but suffer from inferior UA (below 80%).
+These observations underscore the challenges of DM unlearning and highlight the need for further advancements in the field.
+Figure 6: Left: Heatmap visualization of the unlearning accuracy and retainability of ESD on UnlearnCanvas.
 The $x$-axis shows the tested concepts for image generation using the unlearned model, while the $y$-axis indicates the unlearning target.
 Concept types are distinguished by color: styles in blue and objects in orange.
-The figure is separated into different regions to represent corresponding evaluation metrics and the unlearning scopes ($A$ for UA, $B$ for IRA, $C$ for CRA; ‘1’ for style unlearning, ‘2’ for object unlearning).
-Diagonal regions ($A1$ and $A2$) indicate unlearning accuracy, and off-diagonal values ($B1$, $B2$, $C1$, and $C2$) represent retain accuracy.
-Higher values in lighter color denote better performance.
+The figure is divided into regions representing corresponding evaluation metrics and unlearning scopes ($A$ for UA, $B$ for IRA, $C$ for CRA; ‘1’ for style unlearning, ‘2’ for object unlearning).
+Higher values in lighter colors denote better performance.
 The first row serves as a reference for comparison before unlearning.
@@
 Right: Representative cases illustrating each region with images generated before and after unlearning a specific concept.
-6 Experiment Results
-In this section, extensive experiments are conducted on UnlearnCanvas following the evaluation pipeline in Sec.
-5.
-We show that previous MU evaluation, primarily focused on unlearning effectiveness, may present a biased view without considering retainability assessment (Tab.
-3).
-We also find that preserving cross-domain retainability, as quantified by CRA, is more challenging than in-domain retainability (measured by IRA), a previously overlooked aspect in the literature.
-Furthermore, we show that no single MU method consistently excels across all metrics, emphasizing the potential for advancements in existing methods (Fig.
-6).
-In addition, we present a detailed case study on ESD, showcasing how our benchmark offers a comprehensive analysis of an unlearning method’s mechanism on forgetting and retaining information (Fig.
-7).
-Last but not least, we show that different MU methods display distinct unlearning characteristics, which are inherently linked to their underlying unlearning mechanisms (Fig.
-8).
-Overall performance assessment of MU for DMs.
-In Tab.
-3, we provide an overview of the performance of existing unlearning methods, using the metrics discussed in Sec.
-5.
-The key results are summarized below.
-(1) Retainability is essential for comprehensive assessment of MU in DMs.
-Previous evaluations heavily relied on UA (unlearning accuracy), but our assessment reveals a crucial insight: relying solely on UA can provide a skewed view of MU method performance.
-For example, when unlearning artistic painting styles (that we call style unlearning), ESD and UCE achieve similar UA (both around $98\%$), but their retainability (IRA and CRA) differs significantly, with one over $80\%$ and the other around $60\%$.
-Similar conclusions apply to unlearning an image object (that we call object unlearning).
-These findings emphasize the importance of including retainability metrics in the evaluation framework, as demonstrated by UnlearnCanvas.
-(2) CRA is harder to retain than IRA.
-In both style and object unlearning, CRA presents a greater challenge for MU methods compared to IRA.
-This trend is consistent across various methods.
-For example, the UCE method (Gandikota et al., 2023b) exhibits a significant gap between its IRA and CRA, both in style ($60.22\%$ vs.
-$47.71\%$) and object ($39.35\%$ vs.
-$34.67\%$) unlearning scenarios.
-Similar patterns are observed with other methods like FMN and CA.
-This pattern suggests that while MU methods are somewhat effective at preserving concepts within the same domain (in terms of IRA), they face greater challenges in maintaining performance on unlearning target-unrelated concepts across domains (CRA), which has been overlooked in previous studies due to the absence of a multi-label dataset like UnlearnCanvas for systematic unlearning analysis.
-(3) One MU method can perform differently in different domains.
-Take the FMN method (Gandikota et al., 2023b) as an example.
-It exhibits an evident performance difference between the two domains, achieving a UA of $88.48\%$ in style unlearning but a much lower UA of $45.64\%$ in object unlearning.
-In addition, its IRA and CRA metrics show a notable gap, with an IRA of $90.63\%$ for object unlearning compared to $56.77\%$ for style unlearning.
-These examples highlight that a method’s effectiveness in unlearning is not consistent across different scopes.
-(4) No single method can excel in all aspects.
-Each method exhibits unique strengths and weaknesses.
-For example, while ESD (Gandikota et al., 2023a) typically leads an improvement in UA, it may lag behind in terms of IRA and CRA, e.g., in the case of object unlearning.
-Also, the SalUn method (Fan et al., 2023) shows exceptional retainability, with a CRA of $99.59\%$ in object unlearning, but its UA in the same category is lower at $86.91\%$.
-Furthermore, the UCE method (Kumari et al., 2023) enjoys remarkably high UA and efficiency but suffers from degradation in FID and retainability.
-For better visualization, in Fig.
-6, we show this pattern by plotting the performance of different metrics reported in Tab.
-1, illustrating that no method dominates across all radar chart dimensions.
-This underscores the complexity of unlearning in generative models and the need for further advancement.
+A closer look into retainability: A case study on ESD.
+We next perform an in-depth analysis of DMs’ retainability after unlearning, focusing on the most popular ESD method [23].
+In Fig.
+6 (left), we present a heatmap illustrating the per-style/object unlearning accuracy (diagonal) and retain accuracy (off-diagonal) for ESD.
+The heatmap segments are labeled to indicate different regions of interest (ROIs) in style and object unlearning related to our evaluation metrics: A1/A2 for UA, B1/B2 for IRA, and C1/C2 for CRA.
+Key observations are highlighted below.
+First, we observe distinct behaviors of ESD in unlearning styles compared to objects.
+For example, ROI C2 appears darker than C1, suggesting a reduced ability to retain styles during object unlearning.
+Similar patterns are seen between B2 and B1.
+For a clearer explanation, Fig.
+6 (right) provides image generation examples of the DM before and after unlearning within different ROIs.
+Second, We observe that style/object unlearning is relatively easier compared to retaining the generation performance of unlearned DMs conditioned on unlearning-unrelated prompts.
+This is evident in the contrast between ROIs A1/A2 and ROIs B1/B2 or C1/C2.
+One exception is the case of unlearning the object ‘Sea’, which exhibits relatively low UA in A2.
+This may be attributed to post-unlearning image generation still resembling sea waves, as shown by the image examples in Fig.
+6 (right) for A2.
+Third, we observe an inherent trade-off between unlearning effectiveness and retainability.
+For example, while SalUn exhibits more consistent performance across different scenarios than ESD, it does not achieve the same level of UA, as indicated by the lighter diagonal values in Fig.
+A3.
+For the heatmap performance of other unlearning methods, please refer to Appx.
+D.
 Refer to caption Refer to caption Refer to caption
@@
 Figure (c) provides visualizations of generated images using the prompt template ‘A painting of {object} in Sketch style.’ with object being each unlearning target.
-A closer look into retainability: A case study on ESD.
-We next perform a more in-depth analysis of DMs’ retainability after unlearning, with a focus on the ESD method (Gandikota et al., 2023a), which has shown superiority in Tab.
-3 for style unlearning.
-In Fig.
-7 (left), we present a heatmap that illustrates the per-style/object unlearning accuracy (diagonal) and retain accuracy (off-diagonal) for ESD.
-The heatmap segments are labeled to indicate different regions of interest (ROIs) related to our evaluation metrics: A1 and A2 for Unlearning Accuracy (UA) in style and object unlearning, B1 and B2 for In-Domain Retain Accuracy (IRA) in style and object unlearning, and C1 and C2 for Cross-Domain Retain Accuracy (CRA) in style and object unlearning.
-We highlight the following observations.
-First, we observe distinct behaviors of ESD in unlearning styles compared to objects.
-The ROI C2 appears darker than C1, suggesting a reduced ability to retain styles during the unlearning of objects.
-Similar patterns are observable between B2 and B1.
-These findings indicate that an MU method might demonstrate a preference within a specific domain but face challenges in others.
-For a clearer explanation, Fig.
-7 (right) provides image generation examples of the DM before and after unlearning within different ROIs.
-Second, we observe that style/object unlearning is relatively easier compared to retaining the generation performance of unlearned DMs conditioned on unlearning-unrelated prompts.
-This is evident in the contrast between ROIs A1/A2 and ROIs B1/B2 or C1/C2.
-One exception is the case of unlearning the object ‘Sea’, which exhibits relatively low UA in A2.
-This may be attributed to the image generation post-unlearning still resembling sea waves, as shown by the image examples in Fig.
-7 (right) for A2.
-Third, we present a similar accuracy heatmap for SalUn in Fig.
-A1.
-In comparison to ESD, SalUn demonstrates more consistent performance across different unlearning scenarios.
-However, SalUn does not achieve the same level of UA as ESD, as evidenced by the lighter diagonal values in Fig.
-A1.
-This observation underscores the trade-off between unlearning effectiveness and retainability of MU in DMs.
 Understanding unlearning method’s behavior via unlearning directions.
@@
 7 (c), images generated by SalUn post-object unlearning still maintain some object contours (different from the original unlearning target) and better retain style information compared to ESD.
-Expanding applications to style transfer.
-In addition to evaluating DMs’ unlearning performance, it’s worth noting that UnlearnCanvas has the potential to extend its utility to another generative modeling task: style transfer (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021).
-One significant challenge in evaluating style transfer methods is the absence of ground truth images for the given reference style image and content image (Cai et al., 2023).
-Thus, quantitative evaluations are also very limited in the literature.
-In Appx.
-C, we utilize UnlearnCanvas to assess 9 latest style transfer methods, providing enriched and precise evaluations.
-While not the primary focus of this work, this illustrates the great potential of UnlearnCanvas to contribute broader visual generative modeling tasks.
-7 Conclusion
-In this paper, we systematically reviewed the existing the stateful MU (machine unlearning) methods on DMs (diffusion models), and pinpointed the key challenges rooted in their evaluation systems, which could lead to incomplete, inaccurate, and biased assessments.
-In response, we propose UnlearnCanvas, a high-resolution stylized image dataset facilitating to build a comprehensive evaluation system for MU methods, together with a novel systematic retainability measurements, that have been long overlooked but are essential in a all-round evaluation.
-By benchmarking 5 stateful MU methods, we reveal novel insights into the their strengths and weaknesses, deepen the understanding on the underlying mechanism, and further illuminate several promising research directions.
-In the end, with a case study on style transfer, we show UnlearnCanvas enjoys high versatility and great potential to benchmark other generative modeling tasks beyond MU.
-Impact Statement
-This work helps improve the assessment and further promotes the advancement of MU (machine unlearning) methods for DMs (diffusion models), which are known to be effective in relieving or mitigating the various negative societal influences brought by the prevalent usage of DMs, which include but are not limited to the following aspects.
-$\bullet$ Avoiding Copyright Issues.
-There is an urgent need for the generative model providers to scrub the influence of certain data on an already-trained model.
-In January 2023, a notable lawsuit targeted two leading AI art generators, Stable Diffusion (Rombach et al., 2022) and Midjourney (Midjourney et al.,, 2023), for alleged copyright infringement.
-Concurrently, incidents with the recently released Midjourney V6 (Midjourney et al.,, 2023) also highlighted a visual plagiarism issue on famous film scenes.
-These instances illuminate the broad copyright challenges inherent in the way of training data collection method of those foundation generative models’ training datasets.
-MU methods can be used as an effective method to remove the influence of the private data and avoid unnecessary retraining.
-$\bullet$ Mitigating biases and stereotypes.
-Generative AI systems are known to have tendencies towards bias, stereotypes, and reductionism, when it comes to gender, race and national identities (Turk, 2023).
-For example, a recent study on the images generated with Midjourney revealed, that images associated with higher-paying job titles featured people with lighter skin tones, and that results for most professional roles were male-dominated (Vinker et al., 2023).
-MU is known to be effective in eliminating biases rooted in the training data.
-Moreover, UnlearnCanvas offers a flexible framework to benchmark MU techniques against bias removal, allowing for the creation and quantitative control of biases across different object classes for comprehensive bias removal studies.
+Figure 8: UA of DM unlearning against adversarial prompts [59]. Unlearned models in Tab. 2 are used as victim models to generate adversarial prompts.
+4.2 Benchmarking Current DM Unlearning Methods in More Challenging Scenarios
+Unlearning robustness against adversarial prompts.
+Recent studies [59, 60] have highlighted the vulnerabilities of DM unlearning to adversarial prompts, such as jailbreak attacks.
+In Fig.
+8, we use the state-of-the-art attacking method, UnlearnDiffAtk [59], to craft adversarial prompts based on UnlearnCanvas and evaluate the robustness of unlearned DMs shown in Tab.
+2.
+See Appx.
+B.4 for detailed setup.
+As we can see, all the DM unlearning methods experience a significant drop in UA, falling below $60\%$ when confronted with adversarial prompts.
+Notably, methods like UCE, SEOT, and SHS exhibit particularly steep declines.
+Moreover, a high UA under normal conditions does not necessarily imply robustness to adversarial attacks.
+For instance, although UCE achieves a UA over 90% in normal settings, its performance plummets to below 50% against adversarial prompts.
+This stark contrast highlights the importance of the worst-case evaluation for MU methods.
+Table 3: Performance of unlearning style-object combinations.
+The assessment includes UA and retainability in three contexts: SC (style consistency), OC (object consistency), and UP (unrelated prompting).
+The best result in each metric is highlighted in green.
+Method UA SC OC UP
+ESD [23] $91.42\%$ ${4.88\%}$ ${14.72\%}$ $84.38\%$
+FMN [28] ${45.37\%}$ $68.73\%$ $62.74\%$ $83.25\%$
+UCE [24] $75.97\%$ ${4.53\%}$ ${5.72\%}$ ${35.42\%}$
+CA [25] ${47.92\%}$ ${10.08\%}$ ${56.35\%}$ $81.54\%$
+SalUn [27] ${42.21\%}$ $62.45\%$ $70.93\%$ $87.28\%$
+SEOT [30] ${29.32\%}$ ${45.31\%}$ ${53.64\%}$ $85.45\%$
+SPM [26] ${45.72\%}$ ${41.34\%}$ ${36.32\%}$ $67.82\%$
+EDiff [31] $71.33\%$ ${35.23\%}$ ${26.32\%}$ ${51.52\%}$
+SHS [32] ${55.32\%}$ ${14.34\%}$ ${24.32\%}$ $83.95\%$
+Unlearning at a finer scale: Style-object combinations as unlearning targets.
+Next, we consider the fine granularity of the unlearning target, defined by a style-object combination, such as “An image of dogs in Van Gogh style”.
+This unlearning challenge requires the unlearned model to avoid affecting image generation for dogs in non-Van Gogh styles and Van Gogh-style images with non-dog objects.
+See Appx.
+B.5 for detailed setup.
+Here, besides UA, the retainability performance is assessed in three contexts.
+(a) Style consistency (SC): retainability under prompts with different objects but in the same style, e.g., “An image of cats in Van Gogh style”.
+(b) Object consistency (OC): retainability under prompts featuring the same object in different styles, e.g., “An image of dogs in Picasso style”.
+(c) Unrelated prompting (UP): retainability for all other prompts.
+Tab.
+2 presents the performance of unlearning style-object combinations.
+Style-object combinations are more challenging to unlearn than individual objects or styles, as evidenced by a significant drop in UA—over 20% lower compared to values in Tab.
+2.
+Retainability drops to below 20% for top-performing methods like ESD and UCE, originally highlighted for their efficacy.
+This is presumably due to ESD’s underlying unlearning mechanism, which requires only a single prompt, resulting in a poor ability to precisely define the unlearning scope.
+Evaluation in sequential unlearning.
+Furthermore, we evaluate the unlearning performance in sequential unlearning (SU) [70, 71], where unlearning requests arrive sequentially.
+This parallels the continual learning (CL) task, which requires models to not only unlearn new targets effectively but also maintain the unlearning of previous targets while retaining all other knowledge.
+Here, we consider unlearning 6 styles sequentially and the results are presented in Tab.
+A5.
+We remark that the method SEOT does not support sequential unlearning in its original implementation and thus is not included in Tab.
+A5.
+Our findings reveal significant insights.
+(1) Degraded retainability: Sequential unlearning requests generally degrade retainability across all methods, with RA values frequently dropping below the average levels previously seen in Tab.
+2.
+Here RA is given by the average of IRA and CRA.
+(2) Unlearning rebound effect: Knowledge previously unlearned can be inadvertently reactivated by new unlearning requests.
+This is evidenced by decreasing UA values for earlier objectives as more unlearning tasks are introduced.
+This suggests that residual knowledge remains within the model and can be reactivated, aligning with findings from Fig.
+8.
+This indicates the unlearned models by some MU methods do not essentially lose the generation ability of the unlearning target.
+(3) Catastrophic retaining failure: RA significantly drops at a certain request, exemplified by a sudden decrease in RA of UCE from $81.42\%$ to $29.38\%$ after the second request, $\mathcal{T}_{2}$.
+This indicates that the seemingly acceptable side effects generated by some unlearning methods will drastically modify the knowledge representations when accumulated.
+This experiment illuminates the complex dynamics of knowledge removal and retention within DMs and highlights the potential pitfalls of existing unlearning methods when faced with sequential unlearning tasks.
+The observation of the ‘unlearning rebound effect’ and ‘catastrophic retaining failure’ particularly emphasizes the need for a more nuanced understanding of how knowledge is managed within DMs.
+Visualizations.
+We provide plenty of visualizations, illustrations, and qualitative results.
+These visualizations are intended to deepen the understanding of the effects of different MU methods and clearly illustrate the challenges identified in previous sections.
+Specifically, in Fig.
+A11, we provide abundant generation examples of all the 9 methods benchmarked in this work in a case study of unlearning the ‘Cartoon’ style.
+Both the successful and failure cases are demonstrated in the context of unlearning effectiveness, in-domain retainability, and cross-domain retainability.
+Besides, in Fig.
+A12, we provide visualizations for the effect of adversarial prompts enabled by UnlearnDiffAtk [59].
+5 Conclusion, Discussion, and Limitation
+In this paper, we systematically reviewed existing MU (machine unlearning) methods on DMs (diffusion models) and identified key challenges in their evaluation systems that could lead to incomplete, inaccurate, and biased assessments.
+In response, we propose UnlearnCanvas, a high-resolution stylized image dataset designed to facilitate comprehensive evaluation of MU methods.
+We also introduce novel systematic evaluation metrics.
+By benchmarking nine state-of-the-art MU methods, we reveal novel insights into their strengths and weaknesses and also deepen the understanding of their underlying mechanisms.
+Our analysis of three challenging unlearning tasks highlights significant shortcomings of current methods, including a lack of robustness, difficulties in fine-scale unlearning, and issues arising from sequential unlearning tasks.
+These findings also point to meaningful future research directions aimed at developing satisfactory and practical MU methods for real-world applications.
+Furthermore, we recognize limitations in our benchmark, such as its focus on specific Stable Diffusion models and the text-to-image task in DMs.
+We hope this work serves as a foundation for future research to broaden MU evaluations across a wider range of model architectures and tasks in DMs.
+Acknowledgement
+The work of Y.
+Zhang, C.
+Fan, Y.
+Zhang, Y.
+Yao, J.
+Jia, J.
+Liu, and S.
+Liu was supported by the Cisco Research Award, the ARO Award W911NF2310343, the National Science Foundation (NSF) CISE Core Program Award IIS-2207052, and the NSF CAREER Award IIS-2338068, and the Amazon Research Award for AI in Information Security.
+Checklist
+  1. 1.
+    For all authors…
+    1. (a)
+      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+[Yes] The introduction (Sec.
+1) clearly specifies the scope of this paper and the major contributions.
+    2. (b)
+      Did you describe the limitations of your work? [Yes] The limitations are discussed in Sec. 5
+    3. (c)
+      Did you discuss any potential negative societal impacts of your work?
+[No] This work aims to improve the evaluation of machine unlearning technique, which will bring many positive societal impacts, such as mitigating copyright infringement, avoiding sexual or violent content generation, and mitigating the biases and stereotypes of the diffusion models.
+Given the technical focus of this work, there are no specific negative societal consequences directly stemming from it that need to be highlighted here.
+    4. (d)
+      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
+  2. 2.
+    If you are including theoretical results…
+    1. (a)
+      Did you state the full set of assumptions of all theoretical results? [N/A] There are no theoretical results in this paper.
+    2. (b)
+      Did you include complete proofs of all theoretical results? [N/A] There are no theoretical results in this paper.
+  3. 3.
+    If you ran experiments (e.g. for benchmarks)…
+    1. (a)
+      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
+[Yes] There are detailed instructions of the codes in the github repo to facilitate reproducing the results.
+The detailed experiment settings are also disclosed in Appx.
+B.
+    2. (b)
+      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Detailed training settings are disclosed in Sec. 4 and Appx. D.
+    3. (c)
+      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
+[No] The results reported in this work were averaged over dozens of individual unlearning cases, and the variance confirmed to be below 1%.
+We believe this indicates negligible variability.
+Given the large scale of the experiments, error bars were omitted to enhance readability and maintain focus on the key findings.
+    4. (d)
+      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] The compute resources are disclosed in Appx. B.
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
+    1. (a)
+      If your work uses existing assets, did you cite the creators? [Yes] The seed images used to curate UnlearnCanvas belong to the existing assets, and their sources are properly cited in Sec. 3.
+    2. (b)
+      Did you mention the license of the assets? [Yes]
+    3. (c)
+      Did you include any new assets either in the supplemental material or as a URL?
+[Yes] This paper includes a new dataset UnlearnCanvas and the link is released both in the abstract as well as in Appx.
+A.
+    4. (d)
+      Did you discuss whether and how consent was obtained from people whose data you’re using/curating?
+[Yes] The data being used is collected from an open-source photography website and this is mentioned in Sec.
+3.
+    5. (e)
+      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content?
+[Yes] The data does not contain any personally identifiable information or offensive content, as discussed in Appx.
+A.
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects…
+    1. (a)
+      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] No human subjects are involved.
+    2. (b)
+      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] No human subjects are involved.
+    3. (c)
+      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] No human subjects are involved.
 Appendix
-A Detailed Experiment Settings
-A.1 Fine-tuning Style and Object Classifier with UnlearnCanvas.
+[…]
+1
+A UnlearnCanvas Dataset Details
+In this section, we provide a detailed description of the dataset following ‘datasheet for datasets’ [72].
+A.1 Motivation
+This dataset is collected to enable a precise, comprehensive, and automated quantitative evaluation framework for MU (machine unlearning) methods in DMs (diffusion models).
+The current evaluation plans used in the existing literature have exposed several weaknesses, which may lead to incomplete, inaccurate or even biased results, see a more detailed discussion in Sec.
+2.
+To the best of our knowledge, there are no datasets specifically designed to meet the assessing requirements of DM unlearning.
+Therefore, UnlearnCanvas is designed, collected, and made to fill in this gap.
+A.2 Composition: Styles and Object Classes
+There are 60 predetermined artistic styles provided by Fotor [11].
+The images in the same artistic style all share high stylistic consistency, which enables a high-precision style classifier to be trained on them.
+In Fig.
+A9, we list some examples of the images in each style to illustrate these styles.
+There are 20 distinct object classes in UnlearnCanvas.
+In Fig.
+A10, we list some examples of the images in each object to illustrate these classes.
+A.3 Collection and Labeling Process
+The construction of UnlearnCanvas involves two main steps: seed image collection and subsequent image stylization; see Fig.
+3 for an illustration.
+For seed image collection, a set of high-resolution real-world photos are collected from Pexels [69], providing open-sourced photographs.
+There are 20 seed images collected for each of the 20 object classes; see Fig.
+A10.
+After collecting the seed images, we stylize each and every seed image into all 60 predetermined artistic styles; see Fig.
+A9 with Fotor [11].
+After the stylization of all the images, the dataset is structured in a hierarchical manner, and each image is labeled with both its style and object classes.
+In order to support text-to-image training, each image is annotated with the prompt ‘An image of object in style style.’.
+A.4 Uses
+UnlearnCanvas can be used to evaluate MU methods in different unlearning scenarios.
+Please see Sec.
+4 for more details.
+In addition, we stress that UnlearnCanvas can be used for more real-world tasks than unlearning, and we provide an example of how UnlearnCanvas can be used to systematically evaluate another generative task, style transfer, in Appx.
+F.
+We provide very detailed instructions with codes in the GitHub code repository.
+A.5 Distribution
+UnlearnCanvas is an open-sourced dataset and is based on the existing open-sourced data [69].
+We make access to the dataset public under the MIT license.
+We remark that no personally identifiable information or offensive content is included in this dataset.
+The dataset can be accessed either through Google Drive and HuggingFace.
+More resources on the dataset, such as the introduction video and the benchmark, can be found in the official project webpage.
+A.6 Maintenance
+The dataset will be maintained by the lead author Yihua Zhang.
+If needed, the email for contacting is zhan1908@msu.edu.
+The dataset may be updated if needed (with the inclusion of more seed images, more artistic styles, and more objects).
+The updates will be ad-hoc and will not be periodical.
+Each time the dataset is updated, the updates will be reflected in the same GitHub code repository.
+A.7 Author Statements
+The collector and the lead author of this dataset, Yihua Zhang, bears full responsibility for any violation of rights that may arise from the collection of the data included in this research.
+B Reproducibility Statement and Detailed Experiment Settings
+In this section, we provide detailed instructions on the reproduction of our results in Sec.
+4, including the settings of training, the implementation details of the tested machine unlearning methods, and the evaluation details in each unlearning scenario.
+We denote in Sec.
+4, 50 out of the 60 styles proposed in UnlearnCanvas are studied and evaluated, with the other 10 styles being intentionally reserved for future explorations, such as generalizations.
+B.1 Finetuning Style and Object Classifier with UnlearnCanvas
 Style and object classifiers need to be trained as part of the testbed proposed in our evaluation pipeline (Fig.
 5).
-Here, we adopted a pretrained ViT-L/16 model (Dosovitskiy et al., 2020) pretrained on ImageNet and fine-tune it on UnlearnCanvas.
+Here, we adopted a ViT-L/16 model [68] pretrained on ImageNet and finetune it on UnlearnCanvas.
 UnlearnCanvas are split into the train set and test set with a ratio of $9:1$.
-After hyper-parameter tuning, the classifiers are trained with Adam optimizer with a learning rate of $0.01$ for 10 epochs.
-We noticed the model training accuracy can readily converge to $100\%$ even with 10 epochs of training.
-A.2 Fine-tuning Stable Diffusion Model with UnlearnCanvas.
+After hyper-parameter tuning, the classifiers are trained with Adam optimizer at a learning rate of $0.01$ for 10 epochs.
+B.2 Finetuning Stable Diffusion with UnlearnCanvas
 The other part of the testbed is a diffusion model capable of generating high quality images in all the styles associated with all the objects encompassed in UnlearnCanvas in order to guarantee a trustworthy and unbiased evaluation.
 Training settings.
-Practically, we fine-tune the pretrained StableDiffusion v1.5 on UnlearnCanvas for 20k steps with a learning rate of $61e-$ 6.
-Unless otherwise stated, we strictly follow the training configurations used in Stable Diffusion (Rombach et al., 2022).
-For each image in UnlearnCanvas, we annotate the data with text prompt An image of {object} in {style} style, where the object and style are the corresponding object and style label.
+Practically, we finetune the pretrained Stable Diffusion (SD) v1.5 on UnlearnCanvas for 20k steps with a learning rate of $61e-$ 6.
+Unless otherwise stated, we strictly follow the training configurations used in Stable Diffusion [1].
+For each image in UnlearnCanvas, we annotate the data with text prompt An image of {object} in {style}, where the object and style are the corresponding object and style label.
 For seed images, the style label we use is ‘photo’ style.
-We use the training scripts provided by Diffuser official tutorial (StableDiffusion) and the pretraining model card is runwayml/stable-diffusion-v1-5.
+We use the training scripts provided by Diffuser official tutorial (https://huggingface.co/docs/diffusers/v0.13.0/en/training/text2image) and the pretraining model card is runwayml/stable-diffusion-v1-5.
 During training, the checkpoints will be saved every 1000 steps.
 Evaluation.
-To evaluate the quality of the saved checkpoints and selcect the best one for unlearning study, the checkpoints are first used to generate an image set with the same prompt as training (An image of {object} in {style} style) by traversing all the possible style and object labels.
+To evaluate the quality of the saved checkpoints and selcect the best one for unlearning study, the checkpoints are first used to generate an image set with the same prompt as training (An image of {object} in {style}) by traversing all the possible style and object labels.
 Each prompt are used to generate $5$ images with different random seed.
+Each image are sampled with 100 steps with a guidance coeeficient of $9$.
 The image set for each checkpoint are fed into the style and object classifier trained in Appx.
-A.1.
+B.1.
 The model with the highest average performance on all the styles and objects are selected as the testbed for MU study.
@@
 6 (left).
-In Sec.
-6, the experiments were built on the $50$ out of the $60$ styles, which served as the first edition of the dataset, and these styles are listed in Fig.
-7.
-We remark that similar experiments can be done on all the 60 styles.
-A.3 Implementation of the Tested Machine Unlearning Methods
-In this work, we inspected 5 stateful MU methods for DMs. For each method, we use their publicly released source codes as code bases, which are listed below:
+Computing resource.
+In this work, we employ 40×40$\times$40 × NVIDIA RTX A6000 GPUs to conduct all the model training, unlearning, image generation, and evaluations.
+When we finetuned the StableDiffusion on UnlearnCanvas, 8×8$\times$8 × GPUs were used for parallel computing.
+Other experiments were all carried out in a single-GPU environment.
+Around 60,000 GPU hours in total were spent to complete all the experiments.
+B.3 Implementation of DM Unlearning Methods Studied in This Work
+In this work, we inspected a series of stateful MU methods for DMs. For each method, we use their publicly released source codes as code bases, which are listed below:
 In particular, we adopt the following training settings to adapt the methods to our dataset:
-  • ESD: based on the suggestions from the authors, ESD is used to only fine-tune the cross attention-related model weights (ESD-x). Other settings strict follow the ones used in the paper.
+  • ESD: Based on the suggestions from the authors, ESD is used to only finetune the cross attention-related model weights (ESD-x). Other settings strict follow the ones used in the paper.
   • CA: In order to ablate concepts using CA, we first use ChatGPT to generate a list of simple prompts for each concept (including the styles and the objects), namely anchor prompts.
@@
   • UCE: This method requires a guided concept (prompt) for each unlearning concept.
-For style unlearning, we use the prompt An image in {style*} style as the guided concept, where style* represents the next style in UnlearnCanvas in alphabetical order.
+For style unlearning, we use the prompt An image in {style*} as the guided concept, where style* represents the next style in UnlearnCanvas in alphabetical order.
 Similarly, An image of object* is used for object unlearning, where object* is the next object in UnlearnCanvas in alphabetical order.
@@
 For both steps, the mask or model is trained with $10$ epochs.
-B Additional Machine Unlearning Experiment Results
+  • SEOT: To generate unlearned images, we use the prompt An {object*} image in {style*}.
+We then suppress either object* or style* individually.
+Other settings strict follow the ones used in the paper.
+  • SPM: Following the hyperparameters provided by the authors, we trained and obtained Pre-tuned SPMs for all object* and style*.
+During image generation, we combine the pre-tuned SPMs with the DM.
+By calculating the association between words in the prompt and the target word, we determine whether to allow the specified word to preserve, and generate the corresponding image.
+  • EDiff: Based on the authors’ suggestions, EDiff is used to finetune only the cross-attention-related model weights (EraseDiff-x).
+During the unlearning process, we adjusted the hyperparameters, specifically the learning rate and the number of unlearning epochs.
+The model is trained with $5$ epochs.
+  • SHS: SHS consists of two stages: trimming and repairing.
+During the trimming stage, certain weights are re-initialized.
+The repairing stage then restores the model’s utility.
+Throughout the unlearning process, we finetuned the hyperparameters, focusing on the learning rate and the number of unlearning epochs.
+Ultimately, we selected $2$ epochs as the optimal number.
+B.4 Evaluation Details of the Adversarial Prompt Generation for Unlearning Robustness
+In Sec.
+4, we evaluated the robustness of different MU methods against adversarial prompts.
+Here, we use the state-of-the-art method, UnlearnDiffAtk [59] to generate adversarial prompts.
+We set the prepended prompt perturbations by $N=5$ tokens for both style and object unlearning.
+Following the original attack setting in UnlearnDiffAtk [59], to optimize the adversarial perturbations, we sample 50 diffusion time steps and perform PGD running for 40 iterations with a learning rate of 0.01 at each step.
+Prior to projection onto the discrete text space, we utilize the AdamW optimizer.
+B.5 Experiment Details of the Style-Object Combination Unlearning
+Unlearning targets.
+In Sec.
+4, we evaluate the capability of different MU methods on performing unlearning at a finer scale, and we use the style-object combinations as unlearning targets for evaluation.
+Ideally, the UnlearnCanvas dataset can generate 1200 ($60\times 20$) style-object combinations.
+In this work, we randomly select 50 of these combinations for evaluation and we list these combinations below.
+For each method, the same hyper-parameters are used for each MU method as the ones for style and object unlearning in Tab.
+2.
+The unlearning targets include:
+  • ‘An image of Architectures in Abstractionism style.’
+  • ‘An image of Bears in Artist Sketch style.’
+  • ‘An image of Birds in Blossom Season style.’
+  • ‘An image of Butterfly in Bricks style.’
+  • ‘An image of Cats in Byzantine style.’
+  • ‘An image of Dogs in Cartoon style.’
+  • ‘An image of Fishes in Cold Warm style.’
+  • ‘An image of Flame in Color Fantasy style.’
+  • ‘An image of Flowers in Comic Etch style.’
+  • ‘An image of Frogs in Crayon style.’
+  • ‘An image of Horses in Cubism style.’
+  • ‘An image of Human in Dadaism style.’
+  • ‘An image of Jellyfish in Dapple style.’
+  • ‘An image of Rabbits in Defoliation style.’
+  • ‘An image of Sandwiches in Early Autumn style.’
+  • ‘An image of Sea in Expressionism style.’
+  • ‘An image of Statues in Fauvism style.’
+  • ‘An image of Towers in French style.’
+  • ‘An image of Trees in Glowing Sunset style.’
+  • ‘An image of Waterfalls in Gorgeous Love style.’
+  • ‘An image of Architectures in Greenfield style.’
+  • ‘An image of Bears in Impressionism style.’
+  • ‘An image of Birds in Ink Art style.’
+  • ‘An image of Butterfly in Joy style.’
+  • ‘An image of Cats in Liquid Dreams style.’
+  • ‘An image of Dogs in Magic Cube style.’
+  • ‘An image of Fishes in Meta Physics style.’
+  • ‘An image of Flame in Meteor Shower style.’
+  • ‘An image of Flowers in Monet style.’
+  • ‘An image of Frogs in Mosaic style.’
+  • ‘An image of Horses in Neon Lines style.’
+  • ‘An image of Human in On Fire style.’
+  • ‘An image of Jellyfish in Pastel style.’
+  • ‘An image of Rabbits in Pencil Drawing style.’
+  • ‘An image of Sandwiches in Picasso style.’
+  • ‘An image of Sea in Pop Art style.’
+  • ‘An image of Statues in Red Blue Ink style.’
+  • ‘An image of Towers in Rust style.’
+  • ‘An image of Waterfalls in Sketch style.’
+  • ‘An image of Architectures in Sponge Dabbed style.’
+  • ‘An image of Bears in Structuralism style.’
+  • ‘An image of Birds in Superstring style.’
+  • ‘An image of Butterfly in Surrealism style.’
+  • ‘An image of Cats in Ukiyoe style.’
+  • ‘An image of Dogs in Van Gogh style.’
+  • ‘An image of Fishes in Vibrant Flow style.’
+  • ‘An image of Flame in Warm Love style.’
+  • ‘An image of Flowers in Warm Smear style.’
+  • ‘An image of Frogs in Watercolor style.’
+  • ‘An image of Horses in Winter style.’
+Evaluation.
+The evaluation of the style-object combination unlearning concerns four quantitative metrics, one for unlearning effectiveness and three for retainability.
+Before the evaluation, an answer set will be generated exactly following the same procedure introduced in Sec.
+3 and Fig.
+5 after unlearning each target.
+First, the UA (unlearning accuracy) will be evaluated for each answer set, which stands for the ratio of images generated by the target prompt that are neither classified into the target object nor the target style class.
+A high UA denotes a better ability to successfully unlearn the target combination.
+Second, the retainability of generation associated with those prompts close to the unlearning target prompt will be evaluated.
+These prompts can be divided into two groups, the ones sharing the same style but not the object class and the ones sharing the object but not the style class.
+The classification accuracy of the former corresponds to the retainability of the style, i.e., style consistency (SC), while the latter one denotes the object consistency (OC).
+These two quantitative metrics evaluate how well the unlearning method precisely define the unlearning scope and retain the generation ability of those close but innocent concept.
+Thirdly, the retainability of the rest unrelated prompts (UP) are evaluated, which is the last quantitative evaluation metric.
+The results reported in Tab.
+3 are averaged over all the unlearning cases shown above.
+B.6 Experiment Details of Sequential Unlearning
+In Sec.
+4, we also evaluated the MU methods with the task of sequential unlearning (SU), where the efficacy of MU methods in handling multiple sequential unlearning requests $\{\mathcal{T}_{i}\}$ are evaluated.
+This requires models not only to unlearn new targets effectively but also to maintain the unlearning of previous targets, while retaining all other knowledge.
+In the experiments, 6 styles are randomly selected as the unlearning targets and excluded from the RA evaluation.
+The UA of all the already unlearned target will be assessed each time a new request is accomplished.
+The selected 6 styles include:
+  • Abstractionism
+  • Byzantine
+  • Cartoon
+  • Cold Warm
+  • Ukiyoe
+  • Van Gogh
+After each unlearning request, the unlearning effectiveness and retainability are evaluated.
+Specifically, the unlearning accuracy of all the unlearning targets in the previous requests are evaluated to evaluate how the unlearning effect lasts when new unlearning requests arrive.
+In the meantime, the retainability of all the other concepts that are not selected as unlearning targets are evaluated, and to ease the presentation, the retain accuracy of all the concepts (styles and objects) are averaged and reported.
+B.7 Metrics Summary in All Unlearning Settings
+Besides the UnlearnCanvas dataset, the various quantitative evaluation metrics proposed in this work are part of the major contributions to a comprehensive and precise evaluation for DM unlearning methods.
+As there are various unlearning scenarios studied in this work, we provide a summary of these metrics in Tab.
+A1, including their abbreviations, descriptions, and related tables or figures.
+Table A1: A summary of the quantitative metrics used in this work, including their abbreviations, meanings and where they are used.
+Metrics Description Usages (Table & Figure)
+Style/Object Unlearning
+UA Unlearning accuracy Fig. 1, Tab. 1
+IRA In-domain unlearning accuracy Fig. 1, Tab. 1
+CRA Cross-domain unlearning accuracy Fig. 1, Tab. 1
+Unlearning Robustness against Adversarial Prompts
+Rob. Unlearning robustness, unlearning accuracy in the presence of adversarial prompts Fig. 1
+Style-Object Combination Finer-Scale Unlearning
+FU/UA Unlearning accuracy in finer-scale unlearning Fig. 1, Tab. 3
+SC Retainability evaluation of style consistency Tab. 3
+OC Retainability evaluation of object consistency Tab. 3
+UP Retainability evaluation of unrelated prompts Tab. 3
+FR Retainability evaluation in finer-scale unlearning, averaged by SC, OC, and UP Fig. 1
+Sequential or Continual Unlearning
+SU or CU Unlearning accuracy in the context of sequential unlearning Fig. 1, Tab. A5
+SR or CR Retainability in the context of sequential unlearning Fig. 1, Tab. A5
+C A Detailed Comparison between UnlearnCanvas and WikiArt
+Figure A1: Image examples with the same style label from WikiArt [66] and UnlearnCanvas. Images of the same artistic style in UnlearnCanvas exhibit high stylistic consistency compared to WikiArt.
+UnlearnCanvas vs. WikiArt.
+To the best of our knowledge, WikiArt [73] is the most relevant baseline dataset to ours.
+In Tab.
+A2, we provide a direct comparison of the key attributes of these two datasets.
+UnlearnCanvas differs from WikiArt in the following aspects.
+First, UnlearnCanvas includes a greater number of high-resolution images ($15$M) compared to WikiArt ($2$M), a factor that may enhance the training of state-of-the-art DMs.
 Table A2: Comparison with WikiArt, the most relevant dataset containing stylized images to ours.
@@
 Balanced
-WikiArt (Nichol, 2016) $\sim 2M$
+WikiArt [73] $\sim 2M$
 Style-wise
@@
 Object-wise
-Comparison to the most relevant dataset WikiArt.
-We first make a comprehensive comparison between UnlearnCanvas and WikiArt on their key properties in Tab.
-A1.
-As we can see, UnlearnCanvas out performs WikiArt from the several perspectives.
-First, UnlearnCanvas contain images of much higher resolution (15M vs.
-2M).
-This in particular benefits the training associated with DMs, where larger images in general help the model captures more details, such as the stylistic strokes.
-Second, the images in UnlearnCanvas are style-wise supervised.
-This means, for each seed image, a stylized counterpart can be find in each style class.
-This is beneficial for other tasks than MU, such as style transfer, which can provide a ground truth image for precise and robust evaluation.
-This will be detailed in Appx.
-C.
-Third, UnlearnCanvas enjoys high intra-style stylistic consistency and inter-style stylistic distinctiveness.
-This means the images of the same label share much more stylistic similarity, which is easier for a deep learning model to learn and to classify.
-This benefits can also be reflected by the training performance Appx.
-A.1 using UnlearnCanvas and WikiArt.
+Second, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness, as illustrated in Fig.
+A1, where images labeled with ‘Van Gogh Style’ from both datasets are compared.
+In UnlearnCanvas, the images exhibit high stylistic consistency, while WikiArt lacks the necessary clarity for precise assessment.
+This will hamper the MU evaluation as discussed in the challenges (C2) and (C3).
+This benefits can also be reflected by the training performance using UnlearnCanvas and WikiArt.
 The results are reported in Tab.
-2 and Tab.
-A2, respectively.
+A3 and Tab.
+A4, respectively.
 As we can see, the classifier is much more easily trained on UnlearnCanvas, justifying the higher discernible features within each style in UnlearnCanvas.
-Table A2: Style classification results of a ViT-Large (Dosovitskiy et al., 2020) as a style classifier trained on UnlearnCanvas.
+Table A3: Art style reproduction quality using SD v1.5 and SD v2.0 finetuned on WikiArt.
+Images are generated with the prompt “A painting in artist style”, where artist refers to those included in WikiArt.
+The test accuracy on DM-generated images and original WikiArt test images is reported using the style classifier finetuned from the pretrained ViT-L/16 on WikiArt.
+Image
+Source
+Images by
+SD v1.5
+Images by
+SD v2.0
+WikiArt
+Test Set
+Accuracy 41.2% 56.7% 85.4%
+Table A4: Style classification results of a ViT-Large [68] as a style classifier trained on UnlearnCanvas.
 After convergence, the classifier is tested on the test set and the image set generated by SD v1.5 finetuned on UnlearnCanvas.
@@
 Accuracy $100.0\%$ $99.9\%$ $98.8\%$
-A fine-grained comparison per unlearning target: ESD vs. Salun.
-Following the analysis of ESD, we next turn our focus to a comparative analysis with Salun, a method that demonstrated a better balance between unlearning and retaining according to Table 3.
-A similar accuracy heatmap for Salun is presented in Figure A1.
+Third, the images in UnlearnCanvas are style-wise supervised.
+For each seed image, a stylized counterpart can be find in each style class.
+This is beneficial for tasks other than unlearning for text-to-image task, such as image editing, image stylization, and style transfer, which can provide a ground truth image for precise and robust evaluation.
+This will be detailed in Appx.
+F.
+D Additional Experiment Results for Unlearning Evaluation
+D.1 Visualization of the Style and Object Unlearning Performance
+To make a more direct comparison among different MU methods reported in Tab.
+1, the results are visualized in the radar chart Fig.
+A2.
+This figure illustrates that no method dominates across all assessment dimensions.
+This underscores the complexity of unlearning in generative models and the need for further improvement.
+Figure A2: Performance visualization for various unlearning methods as summarized in Table 2.
+For UA, IRA, and CRA, the results are averaged over the style and object unlearning scenarios.
+Other metrics undertake the inverse operation as a smaller values represent better performance.
+Results are normalized to $0\%\sim 100\%$ per metric.
+D.2 A Fine-Grained Comparison per Unlearning Target: ESD vs. Salun
+Following the analysis of ESD in Fig.
+6, we next turn our focus to a comparative analysis with Salun, a method that demonstrated a better balance between unlearning and retaining according to Tab.
+2.
+A similar accuracy heatmap for Salun is presented in Fig.
+A3.
 Compared to ESD, Salun exhibits more consistent performance across various unlearning scenarios, as indicated by the more uniform color distribution in the heatmap.
 This also suggests enhanced retainability.
-However, it is noticeable that Salun does not reach the same level of UA (Unlearning Accuracy) as ESD, as evidenced by the darker diagonal values in Figure A1.
+However, it is noticeable that Salun does not reach the same level of UA (Unlearning Accuracy) as ESD, as evidenced by the darker diagonal values in Fig.
+A3.
 This observation reinforces the existence of a trade-off between unlearning effectiveness and retainability in the visual generative MU task, a phenomenon paralleled in other tasks such as classification.
+D.3 Unlearning Heatmaps of More Unlearning Methods
+[…]
 Figure A3: Heatmap visualization of SalUn. The plot setting is identical to Figure 6.
-C Broader Use Cases of UnlearnCanvas
+Figure A4: Heatmap visualization of FMN. The plot setting is identical to Figure 6.
+Figure A5: Heatmap visualization of SEOT. The plot setting is identical to Figure 6.
+Figure A6: Heatmap visualization of SPM. The plot setting is identical to Figure 6.
+Figure A7: Heatmap visualization of Ediff. The plot setting is identical to Figure 6.
+Figure A8: Heatmap visualization of SHS. The plot setting is identical to Figure 6.
+E Visualizations
+In this section, we aim to provide plenty of visualizations, illustrations, and qualitative results of the dataset and the quantitative results shown in Sec.
+4 and Appx.
+D.
+These visualizations are intended to deepen the understanding of the effects of different MU methods and clearly illustrate the challenges identified in previous sections.
+We hope these visual aids will enable readers to more effectively grasp the nuances of MU methods and their implications for DMs (Diffusion Models).
+Table A5: Performance comparison of different DM unlearning methods in the sequential unlearning setting.
+Each column represents a new unlearning request, denoted by $\mathcal{T}_{i}$, where $\mathcal{T}_{1}$ is the oldest.
+Each row represents the UA for a specific unlearning objective or the retaining accuracy (RA), given by the average of IRA and CRA.
+Results indicating unlearning rebound effect are highlighted in orange, and those signifying catastrophic retaining failure are marked in red.
+Method: ESD Method: FMN
+Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
+————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
+UA $\mathcal{T}_{1}$ $100\%$ $99\%$ $95\%$ $87\%$ $81\%$ $75\%$ UA $\mathcal{T}_{1}$ $88\%$ $99\%$ $99\%$ $98\%$ $99\%$ $99\%$
+$\mathcal{T}_{2}$ $100\%$ $100\%$ $96\%$ $87\%$ $79\%$ $\mathcal{T}_{2}$ $95\%$ $99\%$ $99\%$ $98\%$ $99\%$
+$\mathcal{T}_{3}$ $100\%$ $98\%$ $99\%$ $98\%$ $\mathcal{T}_{3}$ $97\%$ $98\%$ $99\%$ $99\%$
+$\mathcal{T}_{4}$ $100\%$ $100\%$ $99\%$ $\mathcal{T}_{4}$ $99\%$ $99\%$ $99\%$
+$\mathcal{T}_{5}$ $100\%$ $99\%$ $\mathcal{T}_{5}$ $99\%$ $99\%$
+$\mathcal{T}_{6}$ $100\%$ $\mathcal{T}_{6}$ $100\%$
+RA $77.46\%$ $52.94\%$ $35.99\%$ $24.86\%$ $18.69\%$ $12.95\%$ RA $82.39\%$ $14.56\%$ $13.34\%$ $10.42\%$ $9.83\%$ $8.76\%$
+Method: UCE Method: CA
+Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
+————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
+UA $\mathcal{T}_{1}$ $93\%$ $95\%$ $98\%$ $96\%$ $97\%$ $98\%$ UA $\mathcal{T}_{1}$ $58\%$ $55\%$ $59\%$ $45\%$ $44\%$ $40\%$
+$\mathcal{T}_{2}$ $97\%$ $98\%$ $98\%$ $98\%$ $95\%$ $\mathcal{T}_{2}$ $76\%$ $58\%$ $51\%$ $47\%$ $44\%$
+$\mathcal{T}_{3}$ $95\%$ $97\%$ $98\%$ $99\%$ $\mathcal{T}_{3}$ $45\%$ $41\%$ $40\%$ $37\%$
+$\mathcal{T}_{4}$ $98\%$ $98\%$ $98\%$ $\mathcal{T}_{4}$ $71\%$ $70\%$ $60\%$
+$\mathcal{T}_{5}$ $97\%$ $99\%$ $\mathcal{T}_{5}$ $69\%$ $51\%$
+$\mathcal{T}_{6}$ $99\%$ $\mathcal{T}_{6}$ $57\%$
+RA $81.42\%$ $29.38\%$ $18.72\%$ $15.34\%$ $13.32\%$ $11.31\%$ RA $97.24\%$ $93.39\%$ $84.46\%$ $79.32\%$ $71.40\%$ $60.53\%$
+Method: SalUn Method: SPM
+Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
+————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
+UA $\mathcal{T}_{1}$ $84\%$ $79\%$ $78\%$ $65\%$ $67\%$ $64\%$ UA $\mathcal{T}_{1}$ $55\%$ $59\%$ $50\%$ $49\%$ $47\%$ $48\%$
+$\mathcal{T}_{2}$ $81.42\%$ $75\%$ $72\%$ $69\%$ $61\%$ $\mathcal{T}_{2}$ $62\%$ $59\%$ $58\%$ $60\%$ $63\%$
+$\mathcal{T}_{3}$ $90\%$ $85\%$ $84\%$ $87\%$ $\mathcal{T}_{3}$ $42\%$ $39\%$ $40\%$ $41\%$
+$\mathcal{T}_{4}$ $84\%$ $86\%$ $81\%$ $\mathcal{T}_{4}$ $57\%$ $59\%$ $60\%$
+$\mathcal{T}_{5}$ $79\%$ $81\%$ $\mathcal{T}_{5}$ $51\%$ $51\%$
+$\mathcal{T}_{6}$ $89\%$ $\mathcal{T}_{6}$ $43\%$
+RA $85.43\%$ $80.32\%$ $71.42\%$ $65.41\%$ $63.24\%$ $60.19\%$ RA $72.39\%$ $70.42\%$ $67.89\%$ $60.45\%$ $55.32\%$ $51.12\%$
+Method: EDiff Method: SHS
+Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$ Metrics $\mathcal{T}_{1}$ $\mathcal{T}_{1}\sim\mathcal{T}_{2}$ $\mathcal{T}_{1}\sim\mathcal{T}_{3}$ $\mathcal{T}_{1}\sim\mathcal{T}_{4}$ $\mathcal{T}_{1}\sim\mathcal{T}_{5}$ $\mathcal{T}_{1}\sim\mathcal{T}_{6}$
+————————– Unlearning Request ————————– ————————– Unlearning Request ————————–
+UA $\mathcal{T}_{1}$ $97\%$ $93\%$ $91\%$ $93\%$ $85\%$ $90\%$ UA $\mathcal{T}_{1}$ $81\%$ $73\%$ $74\%$ $93\%$ $94\%$ $97\%$
+$\mathcal{T}_{2}$ $92\%$ $89\%$ $93\%$ $91\%$ $87\%$ $\mathcal{T}_{2}$ $69\%$ $61\%$ $89\%$ $94\%$ $97\%$
+$\mathcal{T}_{3}$ $96\%$ $93\%$ $90\%$ $84\%$ $\mathcal{T}_{3}$ $75\%$ $92\%$ $96\%$ $90\%$
+$\mathcal{T}_{4}$ $91\%$ $92\%$ $90.22\%$ $\mathcal{T}_{4}$ $91\%$ $95\%$ $97\%$
+$\mathcal{T}_{5}$ $99\%$ $97\%$ $\mathcal{T}_{5}$ $92\%$ $96\%$
+$\mathcal{T}_{6}$ $94\%$ $\mathcal{T}_{6}$ $94\%$
+RA $92.34\%$ $89.37\%$ $14.35\%$ $12.31\%$ $12.82\%$ $7.42\%$ RA $88.41\%$ $84.32\%$ $73.98\%$ $69.19\%$ $10.76\%$ $10.11\%$
+E.1 Illustrations of the Styles and Objects in UnlearnCanvas
+We first provide an illustration of the styles and object classes included in UnlearnCanvas.
+Specifically, we show the styles in Fig.
+A9 and object classes in Fig.
+A10 with the style and object names disclosed in the captions.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(11) (12) (13) (14) (15) (16) (17) (18) (19) (20)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(21) (22) (23) (24) (25) (26) (27) (28) (29) (30)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(31) (32) (33) (34) (35) (36) (37) (38) (39) (40)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(41) (42) (43) (44) (45) (46) (47) (48) (49) (50)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(51) (52) (53) (54) (55) (56) (57) (58) (59) (60)
+Figure A9: An illustration of the images in each style in UnlearnCanvas used in, which are stylized from the same seed image from the ‘Dogs’ object class.
+The seed image is presented in Fig.
+A10 (6).
+Images are cropped and down-scaled for illustration purpose.
+The name of the styles are: (1) Abstractionism; (2) Artist Sketch; (3) Blossom Season; (4) Blue Blooming; (5) Bricks; (6) Byzantine; (7) Cartoon; (8) Cold Warm; (9) Color Fantasy; (10) Comic Etch; (11) Crayon; (12) Crypto Punks; (13) Cubism; (14) Dadaism; (15) Dapple; (16) Defoliation; (17) Dreamweave; (18) Early Autumn; (19) Expressionism; (20) Fauvism; (21) Foliage Patchwork; (22) French; (23) Glowing Sunset; (24) Gorgeous Love; (25) Greenfield; (26) Impasto; (27) Impressionism; (28) Ink Art; (29) Joy; (30) Liquid Dreams; (31) Palette Knife; (32) Magic Cube; (33) Meta Physics; (34) Meteor Shower; (35) Monet; (36) Mosaic; (37) Neon Lines; (38) On Fire; (39) Pastel; (40) Pencil Drawing; (41) Picasso; (42) Pointillism; (43) Pop Art; (44) Rainwash; (45) Realistic Watercolor; (46) Red Blue Ink; (47) Rust; (48) Sketch; (49) Sponge Dabbed; (50) Structuralism; (51) Superstring; (52) Surrealism; (53) Techno; (54) Ukiyoe; (55) Van Gogh; (56) Vibrant Flow; (57) Warm Love; (58) Warm Smear; (59) Watercolor; (60) Winter.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(11) (12) (13) (14) (15) (16) (17) (18) (19) (20)
+Figure A10: An illustration of the seed images in each object class in UnlearnCanvas.
+Images are cropped and down-scaled for illustration purpose.
+The name of the object classes are: (1) Architecture; (2) Bear; (3) Bird; (4) Butterfly; (5) Cat; (6) Dog; (7) Fish; (8) Flame; (9) Flowers; (10) Frog; (11) Horse; (12) Human; (13) Jellyfish; (14) Rabbits; (15) Sandwich; (16) Sea; (17) Statue; (18) Tower; (19) Tree; (20) Waterfalls.
+E.2 Visualization of Style Unlearning
+In Fig.
+A11, we provide abundant generation examples of all the 9 methods benchmarked in this work in a case study of unlearning the ‘Cartoon’ style.
+Both the successful and failure cases are demonstrated in the context of unlearning effectiveness, in-domain retainability, and cross-domain retainability.
+Figure A11: visualization of the unlearning performance of different methods on the task of style unlearning.
+Three text prompt templates are used to evaluate the unlearning effectiveness, in-domain retainability, and cross-domain retainability of each method.
+Images with green frame denote desirable results, while the ones with red frame denote unlearning or retaining failures.
+E.3 Visualization of the Unlearning Performance in the Presence of Adversarial Prompts
+In Fig.
+A12, we provide visualizations for the effect of adversarial prompts.
+As revealed in Fig.
+8, all the DM unlearning methods experience a significant drop in unlearning effectiveness when attacked by the adversarial prompt, enabled by UnlearnDiffAtk [59].
+We provide the image generation in four unlearning cases (two for style unlearning and two for object unlearning), and show the images of the unlearning target successfully generated in the presence of adversarial prompts.
+Figure A12: Visualization of the images generated by the unlearned DMs using different unlearning methods in the absence or presence of adversarial prompts.
+F Broader Use Cases of UnlearnCanvas
 Although UnlearnCanvas is originally designed for benchmarking MU methods, we would like to demonstrate its broader use cases of benchmarking more generative modeling tasks, thanks to its good properties discussed in Sec.
-4.
+3.
 In this section, we start with a case study on the task of style transfer, which is a much more well-studied topic than MU, but surprisingly also faces great challenges in building up a comprehensive and precise evaluation framework.
@@
 In the end, we will discuss the possibility of using UnlearnCanvas to benchmark more generative modeling tasks.
-C.1 Benchmarking Style Transfer using UnlearnCanvas
+F.1 Benchmarking Style Transfer using UnlearnCanvas
 Figure A13: An illustration of the task of style transfer.
@@
 Figure 1 illustrates the pipeline of this task.
-Most existing literature (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021) utilize WikiArt (Phillips & Mackintosh, 2011) to provide different styles for training the stylization network ${\bm{\theta}}_{s}$.
-During the evaluation, the validation set of WikiArt will serve as the style (reference) images, together with the content images from the COCO dataset (Lin et al., 2014), to form a test bed for style transfer and style learning methods.
+Most existing literature [74, 75, 76, 77, 78, 79] utilize WikiArt [66] to provide different styles for training the stylization network ${\bm{\theta}}_{s}$.
+During the evaluation, the validation set of WikiArt will serve as the style (reference) images, together with the content images from the COCO dataset [80], to form a test bed for style transfer and style learning methods.
 However, such a evaluation scheme has some inherent limitations, which will be detailed below.
 Issues and challenges in the evaluating methods for style transfer
-Although the task of style transfer has been widely studied, its evaluations are still based on very limited quantitative or even sorely qualitative assessments, potentially leading to incomplete and inaccurate assessments (Cai et al., 2023).
+Although the task of style transfer has been widely studied, its evaluations are still based on very limited quantitative or even sorely qualitative assessments, potentially leading to incomplete and inaccurate assessments [81].
 Upon examining the evaluation pipelines of over $10$ state-of-the-art (SOTA) style transfer methods, three significant challenges are identified within the current widely accepted evaluation frameworks.
 $\bullet$ Challenge I (C1): The lack of the ground truth images for style similarity evaluation.
-Unlike other vision tasks like classification (Singh & Singh, 2020), detection (Zou et al., 2023), and segmentation (Minaee et al., 2021), one of the key shortcomings of the current style transfer evaluation lies in the lack of the ground truth images $\mathbf{x}_{g}$ for the given reference style image $\mathbf{x}_{s}$ and the content image $\mathbf{x}_{c}$.
-Consequently, existing evaluation metrics, such as the style loss $\ell_{\mathrm{style}}$ (Gatys et al., 2015a, b), has to be calculated with the reference style image $\mathbf{x}_{s}$ as the ground truth $\mathbf{x}_{g}$, namely $\ell_{\mathrm{style}}(\mathbf{x}_{s},\hat{\mathbf{x}}_{o})$, rather than directly using the ground truth $\ell_{\mathrm{style}}(\mathbf{x}_{o},{\mathbf{x}}_{g})$.
+Unlike other vision tasks like classification [82], detection [83], and segmentation [84], one of the key shortcomings of the current style transfer evaluation lies in the lack of the ground truth images $\mathbf{x}_{g}$ for the given reference style image $\mathbf{x}_{s}$ and the content image $\mathbf{x}_{c}$.
+Consequently, existing evaluation metrics, such as the style loss $\ell_{\mathrm{style}}$ [85, 86], has to be calculated with the reference style image $\mathbf{x}_{s}$ as the ground truth $\mathbf{x}_{g}$, namely $\ell_{\mathrm{style}}(\mathbf{x}_{s},\hat{\mathbf{x}}_{o})$, rather than directly using the ground truth $\ell_{\mathrm{style}}(\mathbf{x}_{o},{\mathbf{x}}_{g})$.
 Obviously, such a indirect evaluation may lead to inaccurate results due to the different contents held in $\mathbf{x}_{s}$ and $\mathbf{x}_{o}$.
-Existing work has demonstrated that such evaluation metrics can lead to very different conclusion from that of the user study (An et al., 2021).
+Existing work has demonstrated that such evaluation metrics can lead to very different conclusion from that of the user study [79].
 Therefore, the creation of a supervised dataset with ground truth stylized images for every content image under each style is a timely remedy.
@@
 Current challenges in such evaluations stem from the lack of reference sets that exhibit high stylistic consistency within each style.
-Figure 4 showcases examples from the widely-used WikiArt dataset.
+Figure A1 showcases examples from the widely-used WikiArt dataset.
 Despite sharing the same artistic label, images within a row show considerable divergence in visual appearance and style.
@@
 To revolutionize the evaluation framework with UnlearnCanvas, it is crucial to first understand the components that form a comprehensive evaluation pipeline, ensuring a thorough and impartial assessment of performance.
-In the realm of style transfer, three pivotal metrics stand paramount, as depicted in Figure,A3:
+In the realm of style transfer, three pivotal metrics stand paramount, as depicted in Figure A14:
 ❶ Style Authenticity: Assesses the extent to which the style of the generated image aligns with that of the provided reference style image(s).
@@
 To encompass the aforementioned aspects of style transfer performance, we propose the following quantitative evaluation metrics:
-① Style Loss: Utilizes a feature map-based style loss (Deng et al., 2022) to quantify the stylistic dissimilarity between image pairs, effectively representing the inverse of style authenticity.
-② Content Loss: Employs a VGG-based, feature-map content loss (Deng et al., 2022; Zhang et al., 2022) to measure the visual dissimilarity between the reference and generated images, essentially mirroring content integrity.
+① Style Loss: Utilizes a feature map-based style loss [87] to quantify the stylistic dissimilarity between image pairs, effectively representing the inverse of style authenticity.
+② Content Loss: Employs a VGG-based, feature-map content loss [87, 88] to measure the visual dissimilarity between the reference and generated images, essentially mirroring content integrity.
 ③ Averaged Standard Deviation (STD): Computes the average STD of style and content loss w.r.t.
@@
 The findings from these comprehensive evaluations are presented in Table A6.
-Following this evaluation pipeline, we scrutinized $9$ prominent style transfer methods, including SaNet (Park & Lee, 2019), Mcc (Deng et al., 2021), Mast (Huo et al., 2021), ArtFlow with its two variants (AF-AdaIn and AF-WCT) (An et al., 2021), IE-Contrast (Chen et al., 2021), Cast (Zhang et al., 2022), StyTr2 (Deng et al., 2022), and BLIP (Li et al., 2023).
+Following this evaluation pipeline, we scrutinized $9$ prominent style transfer methods, including SaNet [78], Mcc [89], Mast [90], ArtFlow with its two variants (AF-AdaIn and AF-WCT) [79], IE-Contrast [91], Cast [88], StyTr2 [87], and BLIP [92].
 Table A6: Performance overview of different style transfer methods evaluated with UnlearnCanvas dataset.
@@
 This is particularly apparent in the style loss evaluation, where Mcc, despite achieving the best result in terms of style loss, exhibits the least stability, denoted by the highest standard deviation.
-C.2 Broader Use Cases of UnlearnCanvas
+F.2 Other Possible Applications of UnlearnCanvas
 In the preceding section, we demonstrated the application of UnlearnCanvas in refining evaluation metrics and frameworks for style transfer.
@@
 Its hierarchical and balanced architecture enables the deliberate introduction of artificial biases by selectively omitting data from specific groups.
-For instance, by predominantly excluding images from styles other than the ‘Van Gogh Style’ within the ‘Dogs’ class, DMs fine-tuned on this dataset will inherently exhibit a tendency to generate images of dogs in the Van Gogh style, particularly when the style is not explicitly specified in the prompt.
+For instance, by predominantly excluding images from styles other than the ‘Van Gogh Style’ within the ‘Dogs’ class, DMs finetuned on this dataset will inherently exhibit a tendency to generate images of dogs in the Van Gogh style, particularly when the style is not explicitly specified in the prompt.
 This approach not only allows for the manipulation and quantification of biases but also paves the way for UnlearnCanvas to become a standardized benchmark for bias mitigation, similar to the role of MU for DMs.
 Vision in-context learning (V-ICL).
-V-ICL (Wang et al., 2023b, a; Bai et al., 2023; Wang et al., 2024) is another domain where UnlearnCanvas can be effectively applied.
+V-ICL [93, 94, 95, 96] is another domain where UnlearnCanvas can be effectively applied.
 The field of V-ICL is in urgent need of robust, comprehensive methods for the fair assessment of existing models.
 In this context, the image pairs from UnlearnCanvas are ideally suited for evaluating various tasks such as style transfer, image inpainting, and image segmentation, offering a rich resource for nuanced and quantitative analyses.
+G Impact Statement
+This work helps improve the assessment and further promotes the advancement of MU (machine unlearning) methods for DMs (diffusion models), which are known to be effective in relieving or mitigating the various negative societal influences brought by the prevalent usage of DMs, which include but are not limited to the following aspects.
+$\bullet$ Avoiding Copyright Issues.
+There is an urgent need for the generative model providers to scrub the influence of certain data on an already-trained model.
+In January 2023, a notable lawsuit targeted two leading AI art generators, Stable Diffusion [1] and Midjourney [2], for alleged copyright infringement.
+Concurrently, incidents with the recently released Midjourney V6 [2] also highlighted a visual plagiarism issue on famous film scenes.
+These instances illuminate the broad copyright challenges inherent in the way of training data collection method of those foundation generative models’ training datasets.
+MU methods can be used as an effective method to remove the influence of the private data and avoid unnecessary retraining.
+$\bullet$ Mitigating biases and stereotypes.
+Generative AI systems are known to have tendencies towards bias, stereotypes, and reductionism, when it comes to gender, race and national identities [17].
+For example, a recent study on the images generated with Midjourney revealed, that images associated with higher-paying job titles featured people with lighter skin tones, and that results for most professional roles were male-dominated [16].
+MU is known to be effective in eliminating biases rooted in the training data.
+Moreover, UnlearnCanvas offers a flexible framework to benchmark MU techniques against bias removal, allowing for the creation and quantitative control of biases across different object classes for comprehensive bias removal studies.